PSPP Developers Guide

GNU PSPP Statistical Analysis Software Release 2.0.1-g1b8d98

Ben Pfaff John Darrington

This manual is for GNU PSPP version 2.0.1-g1b8d98, software for statistical analysis. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1997, 1998, 2004, 2005, 2007, 2010, 2014, 2015, 2016, 2020 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled "GNU Free Documentation License".

Table of Contents

1 System File Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.1 System File Record Structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 File Header Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 Variable Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4 Value Labels Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.5 Document Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.6 Machine Integer Info Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.7 Machine Floating-Point Info Record . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.8 Multiple Response Sets Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.9 Extra Product Info Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 1.10 Variable Display Parameter Record . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 1.11 Variable Sets Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 1.12 Long Variable Names Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 1.13 Very Long String Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 1.14 Character Encoding Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 1.15 Long String Value Labels Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 1.16 Long String Missing Values Record . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 1.17 Data File and Variable Attributes Records . . . . . . . . . . . . . . . . . . . . 22

1.17.1 Variable Roles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 1.18 Extended Number of Cases Record . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 1.19 Other Informational Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 1.20 Dictionary Termination Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 1.21 Data Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

2 SPSS Viewer File Format . . . . . . . . . . . . . . . . . . . . . 28

2.1 Structure Member Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.1.1 The heading Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 2.1.2 The label Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 2.1.3 The container Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.1.4 The text Element (Inside container) . . . . . . . . . . . . . . . . . . . . 34 2.1.5 The html Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.1.6 The table Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.1.7 The graph Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 2.1.8 The model Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 2.1.9 The object and image Elements. . . . . . . . . . . . . . . . . . . . . . . . . . 37 2.1.10 The tree Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 2.1.11 Path Elements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 2.1.12 The pageSetup Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 2.1.13 The text Element (Inside pageParagraph) . . . . . . . . . . . . . . 39

2.2 Light Detail Member Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 2.2.1 Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 2.2.2 Titles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 2.2.3 Footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 2.2.4 Areas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 2.2.5 Borders. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 2.2.6 Print Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 2.2.7 Table Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 2.2.8 Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 2.2.9 Dimensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 2.2.10 Categories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 2.2.11 Axes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 2.2.12 Cells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 2.2.13 Value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 2.2.14 ValueMod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

2.3 Legacy Detail Member Binary Format . . . . . . . . . . . . . . . . . . . . . . . . . 56 2.3.1 Metadata. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 2.3.2 Numeric Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 2.3.3 String Data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

2.4 Legacy Detail Member XML Format . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 2.4.1 The visualization Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 2.4.2 Variable Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

2.4.2.1 The sourceVariable Element . . . . . . . . . . . . . . . . . . . . . . . 62 2.4.2.2 The derivedVariable Element . . . . . . . . . . . . . . . . . . . . . . 63 2.4.2.3 The valueMapEntry Element . . . . . . . . . . . . . . . . . . . . . . . . 64

2.4.3 The extension Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 2.4.4 The graph Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 2.4.5 The location Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 2.4.6 The faceting Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 2.4.7 The facetLayout Element. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 2.4.8 The label Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 2.4.9 The setCellProperties Element. . . . . . . . . . . . . . . . . . . . . . . . . 70 2.4.10 The setFormat Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

2.4.10.1 The numberFormat Element. . . . . . . . . . . . . . . . . . . . . . . . . 73 2.4.10.2 The stringFormat Element. . . . . . . . . . . . . . . . . . . . . . . . . 74 2.4.10.3 The dateTimeFormat Element . . . . . . . . . . . . . . . . . . . . . . 74 2.4.10.4 The elapsedTimeFormat Element. . . . . . . . . . . . . . . . . . . 77 2.4.10.5 The format Element. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 2.4.10.6 The affix Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

2.4.11 The interval Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 2.4.12 The style Element. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 2.4.13 The labelFrame Element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 2.4.14 Legacy Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

3 SPSS TableLook File Formats. . . . . . . . . . . . . . . . . 84

3.1 The .stt Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 3.2 The .tlo Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

3.2.1 PTTableLook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 3.2.2 PVSeparatorStyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 3.2.3 PVCellStyle and PVTextStyle . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

3.2.4 V2Styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

4 Encrypted File Wrappers. . . . . . . . . . . . . . . . . . . . . . 89

4.1 Common Wrapper Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

4.1.1 Checking Passwords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.2 Password Encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

5 Portable File Format. . . . . . . . . . . . . . . . . . . . . . . . . . . 92

5.1 Portable File Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 5.2 Portable File Structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 5.3 Portable File Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 5.4 Version and Date Info Record. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.5 Identification Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.6 Variable Count Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.7 Precision Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.8 Case Weight Variable Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.9 Variable Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.10 Value Label Records. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 5.11 Document Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 5.12 Portable File Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

6 SPSS/PC+ System File Format . . . . . . . . . . . . . . 99

6.1 Record 0: Main Header Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.2 Record 1: Variables Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.3 Record 2: Labels Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 6.4 Record 3: Data Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 6.5 Records 4 and 5: Data Entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

Appendix A GNU Free Documentation License . . 105

1

This manual describes the file formats that PSPP supports.

1 System File Format

An SPSS system file holds a set of cases and dictionary information that describes how they may be interpreted. The system file format dates back 40+ years and has evolved greatly over that time to support new features, but in a way to facilitate interchange between even the oldest and newest versions of software. This chapter describes the system file format.

System files use four data types: 8-bit characters, 32-bit integers, 64-bit integers, and 64-bit floating points, called here char, int32, int64, and flt64, respectively. Data is not necessarily aligned on a word or double-word boundary: the long variable name record (see Section 1.12 [Long Variable Names Record], page 17) and very long string records (see Section 1.13 [Very Long String Record], page 17) have arbitrary byte length and can therefore cause all data coming after them in the file to be misaligned.

Integer data in system files may be big-endian or little-endian. A reader may detect the endianness of a system file by examining layout_code in the file header record (see

[layout_code], page 4).

Floating-point data in system files may nominally be in IEEE 754, IBM, or VAX formats. A reader may detect the floating-point format in use by examining bias in the file header record (see [bias], page 5).

PSPP detects big-endian and little-endian integer formats in system files and translates as necessary. PSPP also detects the floating-point format in use, as well as the endianness of IEEE 754 floating-point numbers, and translates as needed. However, only IEEE 754 numbers with the same endianness as integer data in the same file have actually been observed in system files, and it is likely that other formats are obsolete or were never used.

System files use a few floating point values for special purposes:

SYSMIS The system-missing value is represented by the largest possible negative number in the floating point format (-DBL_MAX).

HIGHEST HIGHEST is used as the high end of a missing value range with an unbounded maximum. It is represented by the largest possible positive number (DBL_MAX).

LOWEST LOWEST is used as the low end of a missing value range with an unbounded minimum. It was originally represented by the second-largest negative number (in IEEE 754 format, 0xffeffffffffffffe). System files written by SPSS 21 and later instead use the largest negative number (-DBL_MAX), the same value as SYSMIS. This does not lead to ambiguity because LOWEST appears in system files only in missing value ranges, which never contain SYSMIS.

System files may use most character encodings based on an 8-bit unit. UTF-16 and UTF-32, based on wider units, appear to be unacceptable. rec_type in the file header record is sufficient to distinguish between ASCII and EBCDIC based encodings. The best way to determine the specific encoding in use is to consult the character encoding record (see Section 1.14 [Character Encoding Record], page 19), if present, and failing that the character_code in the machine integer info record (see Section 1.6 [Machine Integer Info Record], page 10). The same encoding should be used for the dictionary and the data in the file, although it is possible to artificially synthesize files that use different encodings (see Section 1.14 [Character Encoding Record], page 19).

1.1 System File Record Structure

System files are divided into records with the following format:

     int32               type;
     char                data[];

This header does not identify the length of the data or any information about what it contains, so the system file reader must understand the format of data based on type. However, records with type 7, called extension records, have a stricter format:

     int32               type;
     int32               subtype;
     int32               size;
     int32               count;
     char                data[size * count];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. This value identifies a particular kind of extension record.

int32 size;

The size of each piece of data that follows the header, in bytes. Known extension records use 1, 4, or 8, for char, int32, and flt64 format data, respectively.

int32 count;

The number of pieces of data that follow the header.

char data[size * count];

Data, whose format and interpretation depend on the subtype.

An extension record contains exactly size * count bytes of data, which allows a reader that does not understand an extension record to skip it. Extension records provide only nonessential information, so this allows for files written by newer software to preserve backward compatibility with older or less capable readers.

Records in a system file must appear in the following order:

• File header record.

• Variable records.

• All pairs of value labels records and value label variables records, if present.

• Document record, if present.

• Extension (type 7) records, in ascending numerical order of their subtypes. System files written by SPSS include at most one of each kind of extension record. This is generally true of system files written by other software as well, with known exceptions noted below in the individual sections about each type of record.

• Dictionary termination record.

• Data record.

We advise authors of programs that read system files to tolerate format variations. Various kinds of misformatting and corruption have been observed in system files written by SPSS and other software alike. In particular, because extension records provide nonessential information, it is generally better to ignore an extension record entirely than to refuse to read a system file.

The following sections describe the known kinds of records.

1.2 File Header Record

A system file begins with the file header, with the following format:

char                rec_type[4];
char                prod_name[60];
int32               layout_code;
int32               nominal_case_size;
int32               compression;
int32               weight_index;
int32               ncases;
flt64               bias;
char                creation_date[9];
char                creation_time[8];
char                file_label[64];
char                padding[3];
char rec_type[4];

Record type code, either ‘$FL2’ for system files with uncompressed data or data compressed with simple bytecode compression, or ‘$FL3’ for system files with ZLIB compressed data.

This is truly a character field that uses the character encoding as other strings. Thus, in a file with an ASCII-based character encoding this field contains 24 46 4c 32 or 24 46 4c 33, and in a file with an EBCDIC-based encoding this field contains 5b c6 d3 f2. (No EBCDIC-based ZLIB-compressed files have been observed.)

char prod_name[60];

Product identification string. This always begins with the characters ‘@(#) SPSS DATA FILE’. PSPP uses the remaining characters to give its version and the operating system name; for example, ‘GNU pspp 0.1.4 - sparc-sun-solaris2.5.2’. The string is truncated if it would be longer than 60 characters; otherwise it is padded on the right with spaces.

The product name field allow readers to behave differently based on quirks in the way that particular software writes system files. See Section 1.4 [Value Labels Records], page 9, for the detail of the quirk that the PSPP system file reader tolerates in files written by ReadStat, which has

https://github.com/WizardMac/ReadStat in prod_name.
int32 layout_code;

Normally set to 2, although a few system files have been spotted in the wild with a value of 3 here. PSPP use this value to determine the file’s integer endianness (see Chapter 1 [System File Format], page 2).

int32 nominal_case_size;

Number of data elements per case. This is the number of variables, except that long string variables add extra data elements (one for every 8 characters after the first 8). However, string variables do not contribute to this value beyond the first 255 bytes. Further, some software always writes -1 or 0 in this field. In general, it is unsafe for systems reading system files to rely upon this value.

int32 compression;

Set to 0 if the data in the file is not compressed, 1 if the data is compressed with simple bytecode compression, 2 if the data is ZLIB compressed. This field has value 2 if and only if rec_type is ‘$FL3’.

int32 weight_index;

If one of the variables in the data set is used as a weighting variable, set to the dictionary index of that variable, plus 1 (see [Dictionary Index], page 6). Otherwise, set to 0.

int32 ncases;

Set to the number of cases in the file if it is known, or -1 otherwise. In the general case it is not possible to determine the number of cases that will be output to a system file at the time that the header is written. The way that this is dealt with is by writing the entire system file, including the header, then seeking back to the beginning of the file and writing just the ncases field. For files in which this is not valid, the seek operation fails. In this case, ncases remains -1.

flt64 bias;

Compression bias, ordinarily set to 100. Only integers between 1 - bias and 251 - bias can be compressed. By assuming that its value is 100, PSPP uses bias to determine the file’s floating-point format and endianness (see Chapter 1 [System File Format], page 2). If the compression bias is not 100, PSPP cannot auto-detect the floating-point format and assumes that it is IEEE 754 format with the same endianness as the system file’s integers, which is correct for all known system files.

char creation_date[9];

Date of creation of the system file, in ‘dd mmm yy’ format, with the month as standard English abbreviations, using an initial capital letter and following with lowercase. If the date is not available then this field is arbitrarily set to ‘01 Jan

70’.
char creation_time[8];

Time of creation of the system file, in ‘hh:mm:ss’ format and using 24-hour time. If the time is not available then this field is arbitrarily set to ‘00:00:00’.

char file_label[64];

File label declared by the user, if any (see Section “FILE LABEL” in PSPP Users Guide). Padded on the right with spaces. A product that identifies itself as VOXCO INTERVIEWER 4.3 uses CR-only line ends in this field, rather than the more usual LF-only or CR LF line ends.

char padding[3];

Ignored padding bytes to make the structure a multiple of 32 bits in length. Set to zeros.

1.3 Variable Record

There must be one variable record for each numeric variable and each string variable with width 8 bytes or less. String variables wider than 8 bytes have one variable record for each 8 bytes, rounding up. The first variable record for a long string specifies the variable’s correct dictionary information. Subsequent variable records for a long string are filled with dummy information: a type of -1, no variable label or missing values, print and write formats that are ignored, and an empty string as name. A few system files have been encountered that include a variable label on dummy variable records, so readers should take care to parse dummy variable records in the same way as other variable records.

The dictionary index of a variable is a 1-based offset in the set of variable records, including dummy variable records for long string variables. The first variable record has a dictionary index of 1, the second has a dictionary index of 2, and so on.

The system file format does not directly support string variables wider than 255 bytes. Such very long string variables are represented by a number of narrower string variables. See Section 1.13 [Very Long String Record], page 17, for details.

A system file should contain at least one variable and thus at least one variable record, but system files have been observed in the wild without any variables (thus, no data either).

     int32               rec_type;
     int32               type;
     int32               has_var_label;
     int32               n_missing_values;
     int32               print;
     int32               write;
     char                name[8];
     /* Present only if has_var_label is 1. */
     int32               label_len;
     char                label[];

/* Present only if n_missing_values is nonzero. */

     flt64               missing_values[];
int32 rec_type;

Record type code. Always set to 2.

int32 type;

Variable type code. Set to 0 for a numeric variable. For a short string variable or the first part of a long string variable, this is set to the width of the string. For the second and subsequent parts of a long string variable, set to -1, and the remaining fields in the structure are ignored.

int32 has_var_label;

If this variable has a variable label, set to 1; otherwise, set to 0.

int32 n_missing_values;

If the variable has no missing values, set to 0. If the variable has one, two, or three discrete missing values, set to 1, 2, or 3, respectively. If the variable has a range for missing variables, set to -2; if the variable has a range for missing variables plus a single discrete value, set to -3. A long string variable always has the value 0 here. A separate record indicates missing values for long string variables (see Section 1.16 [Long String Missing Values Record], page 21).

int32 print;

Print format for this variable. See below.

int32 write;

Write format for this variable. See below.

char name[8];

Variable name. The variable name must begin with a capital letter or the atsign (‘@’). Subsequent characters may also be digits, octothorpes (‘#’), dollar signs (‘$’), underscores (‘_’), or full stops (‘.’). The variable name is padded on the right with spaces. The ‘name’ fields should be unique within a system file. System files written by SPSS that contain very long string variables with similar names sometimes contain duplicate names that are later eliminated by resolving the very long string names (see Section 1.13 [Very Long String Record], page 17). PSPP handles duplicates by assigning them new, unique names.

int32 label_len;

This field is present only if has_var_label is set to 1. It is set to the length, in characters, of the variable label. The documented maximum length varies from 120 to 255 based on SPSS version, but some files have been seen with longer labels. PSPP accepts labels of any length.

char label[];

This field is present only if has_var_label is set to 1. It has length label_len, rounded up to the nearest multiple of 32 bits. The first label_len characters are the variable’s variable label.

flt64 missing_values[];

This field is present only if n_missing_values is nonzero. It has the same number of 8-byte elements as the absolute value of n_missing_values. Each element is interpreted as a number for numeric variables (with HIGHEST and LOWEST indicated as described in the chapter introduction). For string variables of width less than 8 bytes, elements are right-padded with spaces; for string variables wider than 8 bytes, only the first 8 bytes of each missing value are specified, with the remainder implicitly all spaces. For discrete missing values, each element represents one missing value. When a range is present, the first element denotes the minimum value in the range, and the second element denotes the maximum value in the range. When a range plus a value are present, the third element denotes the additional discrete missing value.

The print and write members of sysfile variable are output formats coded into int32 types. The least-significant byte of the int32 represents the number of decimal places, and the next two bytes in order of increasing significance represent field width and format type, respectively. The most-significant byte is not used and should be set to zero.

Format types are defined as follows:

Value Meaning
0  Not used.

1  A
2  AHEX
3  COMMA
4  DOLLAR
5  F
6  IB
7  PIBHEX
8  P
9  PIB
10  PK
11  RB
12  RBHEX

13  Not used.
14  Not used.
15  Z
16  N
17  E
18  Not used.
19  Not used.

20  DATE
21  TIME
22  DATETIME
23  ADATE
24  JDATE
25  DTIME
26  WKDAY
27  MONTH
28  MOYR
29  QYR
30  WKYR
31  PCT
32  DOT
33  CCA
34  CCB
35  CCC
36  CCD
37  CCE
38  EDATE
39  SDATE

     40  MTIME
     41  YMDHMS

A few system files have been observed in the wild with invalid write fields, in particular with value 0. Readers should probably treat invalid print or write fields as some default format.

1.4 Value Labels Records

The value label records documented in this section are used for numeric and short string variables only. Long string variables may have value labels, but their value labels are recorded using a different record type (see Section 1.15 [Long String Value Labels Record], page 19).

ReadStat (see Section 1.2 [File Header Record], page 4) writes value labels that label a single value more than once. In more detail, it emits value labels whose values are longer than string variables’ widths, that are identical in the actual width of the variable, e.g. labels for values ABC123 and ABC456 for a string variable with width 3. For files written by this software, PSPP ignores such labels.

The value label record has the following format:

     int32               rec_type;
     int32               label_count;

/* Repeated n_label times. */

     char                value[8];
     char                label_len;
     char                label[];
int32 rec_type;

Record type. Always set to 3.

int32 label_count;

Number of value labels present in this record.

The remaining fields are repeated count times. Each repetition specifies one value label.

char value[8];

A numeric value or a short string value padded as necessary to 8 bytes in length. Its type and width cannot be determined until the following value label variables record (see below) is read.

char label_len;

The label’s length, in bytes. The documented maximum length varies from 60 to 120 based on SPSS version. PSPP supports value labels up to 255 bytes long.

char label[];

label_len bytes of the actual label, followed by up to 7 bytes of padding to bring label and label_len together to a multiple of 8 bytes in length.

The value label record is always immediately followed by a value label variables record with the following format:

     int32               rec_type;

int32               var_count;
int32               vars[];
int32 rec_type;

Record type. Always set to 4.

int32 var_count;

Number of variables that the associated value labels from the value label record are to be applied.

int32 vars[];

A list of 1-based dictionary indexes of variables to which to apply the value labels (see [Dictionary Index], page 6). There are var_count elements. String variables wider than 8 bytes may not be specified in this list.

1.5 Document Record

The document record, if present, has the following format:

int32               rec_type;
int32               n_lines;
char                lines[][80];
int32 rec_type;

Record type. Always set to 6.

int32 n_lines;

Number of lines of documents present. This should be greater than zero, but ReadStats writes system files with zero n_lines.

char lines[][80];

Document lines. The number of elements is defined by n_lines. Lines shorter than 80 characters are padded on the right with spaces.

1.6 Machine Integer Info Record

The integer info record, if present, has the following format:

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Data. */

int32               version_major;
int32               version_minor;
int32               version_revision;
int32               machine_code;
int32               floating_point_rep;
int32               compression_code;
int32               endianness;

int32               character_code;
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 3.

int32 size;

Size of each piece of data in the data part, in bytes. Always set to 4.

int32 count;

Number of pieces of data in the data part. Always set to 8.

int32 version_major;

PSPP major version number. In version x.y.z, this is x.

int32 version_minor;

PSPP minor version number. In version x.y.z, this is y.

int32 version_revision;

PSPP version revision number. In version x.y.z, this is z.

int32 machine_code;

Machine code. PSPP always set this field to value to -1, but other values may appear.

int32 floating_point_rep;

Floating point representation code. For IEEE 754 systems this is 1. IBM 370 sets this to 2, and DEC VAX E to 3.

int32 compression_code;

Compression code. Always set to 1, regardless of whether or how the file is compressed.

int32 endianness;

Machine endianness. 1 indicates big-endian, 2 indicates little-endian.

int32 character_code;

Character code. The following values have been actually observed in system files: 1 EBCDIC. 2 7-bit ASCII. 1250 The windows-1250 code page for Central European and Eastern

European languages.

1252 The windows-1252 code page for Western European languages. 28591 ISO 8859-1. 65001 UTF-8. The following additional values are known to be defined: 3 8-bit “ASCII”.

4 DEC Kanji.

Other Windows code page numbers are known to be generally valid. Old versions of SPSS for Unix and Windows always wrote value 2 in this field, regardless of the encoding in use. Newer versions also write the character encoding as a string (see Section 1.14 [Character Encoding Record], page 19).

1.7 Machine Floating-Point Info Record

The floating-point info record, if present, has the following format:

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Data. */

flt64               sysmis;
flt64               highest;
flt64               lowest;
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 4.

int32 size;

Size of each piece of data in the data part, in bytes. Always set to 8.

int32 count;

Number of pieces of data in the data part. Always set to 3.

flt64 sysmis;
flt64 highest;
flt64 lowest;

The system missing value, the value used for HIGHEST in missing values, and the value used for LOWEST in missing values, respectively. See Chapter 1 [System File Format], page 2, for more information. The SPSSWriter library in PHP, which identifies itself as FOM SPSS 1.0.0 in the file header record prod_name field, writes unexpected values to these fields, but it uses the same values consistently throughout the rest of the file.

1.8 Multiple Response Sets Records

The system file format has two different types of records that represent multiple response sets (see Section “MRSETS” in PSPP Users Guide). The first type of record describes multiple response sets that can be understood by SPSS before version 14. The second type of record, with a closely related format, is used for multiple dichotomy sets that use the CATEGORYLABELS=COUNTEDVALUES feature added in version 14.

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Exactly count bytes of data. */

char                mrsets[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Set to 7 for records that describe multiple response sets understood by SPSS before version 14, or to 19 for records that describe dichotomy sets that use the CATEGORYLABELS=COUNTEDVALUES feature added in version 14.

int32 size;

The size of each element in the mrsets member. Always set to 1.

int32 count;

The total number of bytes in mrsets.

char mrsets[];

Zero or more line feeds (byte 0x0a), followed by a series of multiple response sets, each of which consists of the following:

• The set’s name (an identifier that begins with ‘$’), in mixed upper and lower case.

• An equals sign (‘=’).

• ‘C’ for a multiple category set, ‘D’ for a multiple dichotomy set with CAT-EGORYLABELS=VARLABELS, or ‘E’ for a multiple dichotomy set with CATEGORYLABELS=COUNTEDVALUES.

• For a multiple dichotomy set with CATEGORYLABELS=COUNTEDVALUES, a space, followed by a number expressed as decimal digits, followed by a space. If LABELSOURCE=VARLABEL was specified on MRSETS, then the number is 11; otherwise it is 1.^{1}

• For either kind of multiple dichotomy set, the counted value, as a positive integer count specified as decimal digits, followed by a space, followed by as many string bytes as specified in the count. If the set contains numeric variables, the string consists of the counted integer value expressed as decimal digits. If the set contains string variables, the string contains the counted string value. Either way, the string may be padded on the right with spaces (older versions of SPSS seem to always pad to a width of 8 bytes; newer versions don’t).

• A space.

1 This part of the format may not be fully understood, because only a single example of each possibility

has been examined.

• The multiple response set’s label, using the same format as for the counted value for multiple dichotomy sets. A string of length 0 means that the set does not have a label. A string of length 0 is also written if LABEL-SOURCE=VARLABEL was specified.

• A space.

• The short names of the variables in the set, converted to lowercase, each separated from the previous by a single space. Even though a multiple response set must have at least two variables, some system files contain multiple response sets with no variables or one variable. The source and meaning of these multiple response sets is unknown. (Perhaps they arise from creating a multiple response set then deleting all the variables that it contains?)

• One line feed (byte 0x0a). Sometimes multiple, even hundreds, of line feeds are present.

Example: Given appropriate variable definitions, consider the following MRSETS command:

MRSETS /MCGROUP NAME=$a LABEL='my mcgroup' VARIABLES=a b c
       /MDGROUP NAME=$b VARIABLES=g e f d VALUE=55
       /MDGROUP NAME=$c LABEL='mdgroup #2' VARIABLES=h i j VALUE='Yes'
       /MDGROUP NAME=$d LABEL='third mdgroup' CATEGORYLABELS=COUNTEDVALUES
        VARIABLES=k l m VALUE=34
       /MDGROUP NAME=$e CATEGORYLABELS=COUNTEDVALUES LABELSOURCE=VARLABEL
        VARIABLES=n o p VALUE='choice'.

The above would generate the following multiple response set record of subtype 7:

$a=C 10 my mcgroup a b c
$b=D2 55 0  g e f d
$c=D3 Yes 10 mdgroup #2 h i j

It would also generate the following multiple response set record with subtype 19:

$d=E 1 2 34 13 third mdgroup k l m
$e=E 11 6 choice 0  n o p

1.9 Extra Product Info Record

This optional record appears to contain a text string that describes the program that wrote the file and the source of the data. (This is redundant with the file label and product info found in the file header record.)

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Exactly count bytes of data. */

     char                info[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 10.

int32 size;

The size of each element in the info member. Always set to 1.

int32 count;

The total number of bytes in info.

char info[];

A text string. A product that identifies itself as VOXCO INTERVIEWER 4.3 uses CR-only line ends in this field, rather than the more usual LF-only or CR LF line ends.

1.10 Variable Display Parameter Record

The variable display parameter record, if present, has the following format:

/* Header. */

     int32               rec_type;
     int32               subtype;
     int32               size;
     int32               count;

/* Repeated count times. */

     int32               measure;

int32 width; /* Not always present. */

     int32               alignment;
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 11.

int32 size;

The size of int32. Always set to 4.

int32 count;

The number of sets of variable display parameters (ordinarily the number of variables in the dictionary), times 2 or 3.

The remaining members are repeated count times, in the same order as the variable records. No element corresponds to variable records that continue long string variables. The meanings of these members are as follows:

int32 measure;

The measurement level of the variable:

0 Unknown 1 Nominal

2 Ordinal

3 Scale

An “unknown” measure of 0 means that the variable was created in some way that doesn’t make the measurement level clear, e.g. with a COMPUTE transformation. PSPP sets the measurement level the first time it reads the data using the rules documented in Section “Measurement Level” in PSPP Users Guide, so this should rarely appear.

int32 width;

The width of the display column for the variable in characters. This field is present if count is 3 times the number of variables in the dictionary. It is omitted if count is 2 times the number of variables.

int32 alignment;

The alignment of the variable for display purposes:

0 Left aligned

1 Right aligned

2 Centre aligned

1.11 Variable Sets Record

The SPSS GUI offers users the ability to arrange variables in sets. Users may enable and disable sets individually, and the data editor and analysis dialog boxes only show enabled sets. Syntax does not use variable sets.

The variable sets record, if present, has the following format: /* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Exactly count bytes of text. */

char                text[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 5.

int32 size;

Always set to 1.

int32 count;

The total number of bytes in text.

char text[];

The variable sets, in a text-based format. Each variable set occupies one line of text, each of which ends with a line feed (byte 0x0a), optionally preceded by a carriage return (byte 0x0d). Each line begins with the name of the variable set, followed by an equals sign (‘=’) and a space (byte 0x20), followed by the long variable names of the members of the set, separated by spaces. A variable set may be empty, in which case the equals sign and the space following it are still present.

1.12 Long Variable Names Record

If present, the long variable names record has the following format:

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Exactly count bytes of data. */

char                var_name_pairs[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 13.

int32 size;

The size of each element in the var_name_pairs member. Always set to 1.

int32 count;

The total number of bytes in var_name_pairs.

char var_name_pairs[];

A list of key–value tuples, where key is the name of a variable, and value is its long variable name. The key field is at most 8 bytes long and must match the name of a variable which appears in the variable record (see Section 1.3 [Variable Record], page 6). The value field is at most 64 bytes long. The key and value fields are separated by a ‘=’ byte. Each tuple is separated by a byte whose value is 09. There is no trailing separator following the last tuple. The total length is count bytes.

1.13 Very Long String Record

Old versions of SPSS limited string variables to a width of 255 bytes. For backward compatibility with these older versions, the system file format represents a string longer than 255 bytes, called a very long string, as a collection of strings no longer than 255 bytes each. The strings concatenated to make a very long string are called its segments; for consistency, variables other than very long strings are considered to have a single segment.

A very long string with a width of w has n = (w + 251) / 252 segments, that is, one segment for every 252 bytes of width, rounding up. It would be logical, then, for each of the segments except the last to have a width of 252 and the last segment to have the remainder, but this is not the case. In fact, each segment except the last has a width of 255 bytes. The last segment has width w - (n - 1) * 252; some versions of SPSS make it slightly wider, but not wide enough to make the last segment require another 8 bytes of data.

Data is packed tightly into segments of a very long string, 255 bytes per segment. Because 255 bytes of segment data are allocated for every 252 bytes of the very long string’s width (approximately), some unused space is left over at the end of the allocated segments. Data in unused space is ignored.

Example: Consider a very long string of width 20,000. Such a very long string has 20,000 / 252 = 80 (rounding up) segments. The first 79 segments have width 255; the last segment has width 20,000 - 79 * 252 = 92 or slightly wider (up to 96 bytes, the next multiple of 8). The very long string’s data is actually stored in the 19,890 bytes in the first 78 segments, plus the first 110 bytes of the 79th segment (19,890 + 110 = 20,000). The remaining 145 bytes of the 79th segment and all 92 bytes of the 80th segment are unused.

The very long string record explains how to stitch together segments to obtain very long string data. For each of the very long string variables in the dictionary, it specifies the name of its first segment’s variable and the very long string variable’s actual width. The remaining segments immediately follow the named variable in the system file’s dictionary.

The very long string record, which is present only if the system file contains very long string variables, has the following format:

/* Header. */

     int32               rec_type;
     int32               subtype;
     int32               size;
     int32               count;

/* Exactly count bytes of data. */

     char                string_lengths[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 14.

int32 size;

The size of each element in the string_lengths member. Always set to 1.

int32 count;

The total number of bytes in string_lengths.

char string_lengths[];

A list of key–value tuples, where key is the name of a variable, and value is its length. The key field is at most 8 bytes long and must match the name of a variable which appears in the variable record (see Section 1.3 [Variable Record], page 6). The value field is exactly 5 bytes long. It is a zero-padded, ASCII-encoded string that is the length of the variable. The key and value fields are separated by a ‘=’ byte. Tuples are delimited by a two-byte sequence {00, 09}. After the last tuple, there may be a single byte 00, or {00, 09}. The total length is count bytes.

1.14 Character Encoding Record

This record, if present, indicates the character encoding for string data, long variable names, variable labels, value labels and other strings in the file.

/* Header. */

     int32               rec_type;
     int32               subtype;
     int32               size;
     int32               count;

/* Exactly count bytes of data. */

     char                encoding[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 20.

int32 size;

The size of each element in the encoding member. Always set to 1.

int32 count;

The total number of bytes in encoding.

char encoding[];

The name of the character encoding. Normally this will be an official IANA

          character set name or alias. See http://www.iana.org/assignments/

character-sets. Character set names are not case-sensitive, but SPSS appears to write them in all-uppercase.

This record is not present in files generated by older software. See also the character_ code field in the machine integer info record (see [character-code], page 11).

When the character encoding record and the machine integer info record are both present, all system files observed in practice indicate the same character encoding, e.g. 1252 as

character_code and windows-1252 as encoding, 65001 and UTF-8, etc.

If, for testing purposes, a file is crafted with different character_code and encoding, it seems that character_code controls the encoding for all strings in the system file before the dictionary termination record, including strings in data (e.g. string missing values), and encoding controls the encoding for strings following the dictionary termination record.

1.15 Long String Value Labels Record

This record, if present, specifies value labels for long string variables.

/* Header. */

     int32               rec_type;

int32               subtype;
int32               size;
int32               count;

/* Repeated up to exactly count bytes. */

int32               var_name_len;
char                var_name[];
int32               var_width;
int32               n_labels;
long_string_label   labels[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 21.

int32 size;

Always set to 1.

int32 count;

The number of bytes following the header until the next header.

int32 var_name_len;
char var_name[];

The number of bytes in the name of the variable that has long string value labels, plus the variable name itself, which consists of exactly var_name_len bytes. The variable name is not padded to any particular boundary, nor is it null-terminated.

int32 var_width;

The width of the variable, in bytes, which will be between 9 and 32767.

int32 n_labels;
long_string_label labels[];

The long string labels themselves. The labels array contains exactly n_labels elements, each of which has the following substructure:

     int32               value_len;
     char                value[];
     int32               label_len;
     char                label[];
int32 value_len;
char value[];

The string value being labeled. value_len is the number of bytes in value; it is equal to var_width. The value array is not padded or null-terminated.

int32 label_len;
char label[];

The label for the string value. label_len, which must be between 0 and 120, is the number of bytes in label. The label array is not padded or null-terminated.

1.16 Long String Missing Values Record

This record, if present, specifies missing values for long string variables.

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Repeated up to exactly count bytes. */

int32               var_name_len;
char                var_name[];
char                n_missing_values;
int32               value_len;
char                values[values_len * n_missing_values];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 22.

int32 size;

Always set to 1.

int32 count;

The number of bytes following the header until the next header.

int32 var_name_len;
char var_name[];

The number of bytes in the name of the long string variable that has missing values, plus the variable name itself, which consists of exactly var_name_len bytes. The variable name is not padded to any particular boundary, nor is it null-terminated.

char n_missing_values;

The number of missing values, either 1, 2, or 3. (This is, unusually, a single byte instead of a 32-bit number.)

int32 value_len;

The length of each missing value string, in bytes. This value should be 8, because long string variables are at least 8 bytes wide (by definition), only the first 8 bytes of a long string variable’s missing values are allowed to be nonspaces, and any spaces within the first 8 bytes are included in the missing value here.

char values[values_len * n_missing_values]

The missing values themselves, without any padding or null terminators.

An earlier version of this document stated that value_len was repeated before each of the missing values, so that there was an extra int32 value of 8 before each missing value after the first. Old versions of PSPP wrote data files in this format. Readers can tolerate this mistake, if they wish, by noticing and skipping the extra int32 values, which wouldn’t ordinarily occur in strings.

1.17 Data File and Variable Attributes Records

The data file and variable attributes records represent custom attributes for the system file or for individual variables in the system file, as defined on the DATAFILE ATTRIBUTE (see Section “DATAFILE ATTRIBUTE” in PSPP Users Guide) and VARIABLE ATTRIBUTE commands (see Section “VARIABLE ATTRIBUTE” in PSPP Users Guide), respectively.

/* Header. */

int32               rec_type;
int32               subtype;
int32               size;
int32               count;

/* Exactly count bytes of data. */

char                attributes[];
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 17 for a data file attribute record or to 18 for a variable attributes record.

int32 size;

The size of each element in the attributes member. Always set to 1.

int32 count;

The total number of bytes in attributes.

char attributes[];

The attributes, in a text-based format. In record subtype 17, this field contains a single attribute set. An attribute set is a sequence of one or more attributes concatenated together. Each attribute consists of a name, which has the same syntax as a variable name, followed by, inside parentheses, a sequence of one or more values. Each value consists of a string enclosed in single quotes (') followed by a line feed (byte 0x0a). A value may contain single quote characters, which are not themselves escaped or quoted or required to be present in pairs. There is no apparent way to embed a line feed in a value. There is no distinction between an attribute with a single value and an attribute array with one element. In record subtype 18, this field contains a sequence of one or more variable attribute sets. If more than one variable attribute set is present, each one after the first is delimited from the previous by /. Each variable attribute set consists of a long variable name, followed by :, followed by an attribute set with the same syntax as on record subtype 17.

System files written by Stata 14.1/-savespss- 1.77 by S.Radyakin may in-

clude multiple records with subtype 18, one per variable that has variable attributes.

The total length is count bytes.

Example

A system file produced with the following VARIABLE ATTRIBUTE commands in effect:

     VARIABLE ATTRIBUTE VARIABLES=dummy ATTRIBUTE=fred[1]('23') fred[2]('34').
     VARIABLE ATTRIBUTE VARIABLES=dummy ATTRIBUTE=bert('123').

will contain a variable attribute record with the following contents:

     0000  07 00 00 00 12 00 00 00  01 00 00 00 22 00 00 00  |............"...|
     0010  64 75 6d 6d 79 3a 66 72  65 64 28 27 32 33 27 0a  |dummy:fred('23'.|
     0020  27 33 34 27 0a 29 62 65  72 74 28 27 31 32 33 27  |'34'.)bert('123'|
     0030  0a 29                                             |.)              |

1.17.1 Variable Roles

A variable’s role is represented as an attribute named $@Role. This attribute has a single element whose values and their meanings are:

0 Input. This, the default, is the most common role.

1 Output.

2 Both.

3 None.

4 Partition.

5 Split.

1.18 Extended Number of Cases Record

The file header record expresses the number of cases in the system file as an int32 (see Section 1.2 [File Header Record], page 4). This record allows the number of cases in the system file to be expressed as a 64-bit number.

     int32               rec_type;
     int32               subtype;
     int32               size;
     int32               count;
     int64               unknown;
     int64               ncases64;
int32 rec_type;

Record type. Always set to 7.

int32 subtype;

Record subtype. Always set to 16.

int32 size;

Size of each element. Always set to 8.

int32 count;

Number of pieces of data in the data part. Alway set to 2.

int64 unknown;

Meaning unknown. Always set to 1.

int64 ncases64;

Number of cases in the file as a 64-bit integer. Presumably this could be -1 to indicate that the number of cases is unknown, for the same reason as ncases in the file header record, but this has not been observed in the wild.

1.19 Other Informational Records

This chapter documents many specific types of extension records are documented here, but others are known to exist. PSPP ignores unknown extension records when reading system files.

The following extension record subtypes have also been observed, with the following believed meanings:

6 Date info, probably related to USE (according to Aapi Hämäläinen).

12 A UUID in the format described in RFC 4122. Only two examples observed, both written by SPSS 13, and in each case the UUID contained both upper and lower case.

24 XML that describes how data in the file should be displayed on-screen.

1.20 Dictionary Termination Record

The dictionary termination record separates all other records from the data records.

     int32               rec_type;
     int32               filler;
int32 rec_type;

Record type. Always set to 999.

int32 filler;

Ignored padding. Should be set to 0.

1.21 Data Record

The data record must follow all other records in the system file. Every system file must have a data record that specifies data for at least one case. The format of the data record varies depending on the value of compression in the file header record:

0: no compression Data is arranged as a series of 8-byte elements. Each element corresponds to the variable declared in the respective variable record (see Section 1.3 [Variable Record], page 6). Numeric values are given in flt64 format; string values are literal characters string, padded on the right when necessary to fill out 8-byte units.

1: bytecode compression The first 8 bytes of the data record is divided into a series of 1-byte command codes. These codes have meanings as described below:

0 Ignored. If the program writing the system file accumulates compressed data in blocks of fixed length, 0 bytes can be used to pad out extra bytes remaining at the end of a fixed-size block.

1 through 251 A number with value code - bias, where code is the value of the compression code and bias is the variable bias from the file header. For example, code 105 with bias 100.0 (the normal value) indicates a numeric variable of value 5. A code of 0 (after subtracting the bias) in a string field encodes null bytes. This is unusual, since a string field normally encodes text data, but it exists in real system files.

252 End of file. This code may or may not appear at the end of the data stream. PSPP always outputs this code but its use is not required.

253 A numeric or string value that is not compressible. The value is stored in the 8 bytes following the current block of command bytes. If this value appears twice in a block of command bytes, then it indicates the second group of 8 bytes following the command bytes, and so on.

254 An 8-byte string value that is all spaces.

255 The system-missing value.

The end of the 8-byte group of bytecodes is followed by any 8-byte blocks of non-compressible values indicated by code 253. After that follows another 8byte group of bytecodes, then those bytecodes’ non-compressible values. The pattern repeats to the end of the file or a code with value 252.

2: ZLIB compression

The data record consists of the following, in order:

• ZLIB data header, 24 bytes long.

• One or more variable-length blocks of ZLIB compressed data.

• ZLIB data trailer, with a 24-byte fixed header plus an additional 24 bytes for each preceding ZLIB compressed data block.

The ZLIB data header has the following format:

int64               zheader_ofs;
int64               ztrailer_ofs;
int64               ztrailer_len;
int64 zheader_ofs;

The offset, in bytes, of the beginning of this structure within the system file.

int64 ztrailer_ofs;

The offset, in bytes, of the first byte of the ZLIB data trailer.

int64 ztrailer_len;

The number of bytes in the ZLIB data trailer. This and the previous field sum to the size of the system file in bytes.

The data header is followed by (ztrailer_len - 24) / 24 ZLIB compressed data blocks. Each ZLIB compressed data block begins with a ZLIB header as specified in RFC 1950, e.g. hex bytes 78 01 (the only header yet observed in practice). Each block decompresses to a fixed number of bytes (in practice only 0x3ff000-byte blocks have been observed), except that the last block of data may be shorter. The last ZLIB compressed data block gends just before offset

ztrailer_ofs.

The result of ZLIB decompression is bytecode compressed data as described above for compression format 1. The ZLIB data trailer begins with the following 24-byte fixed header:

     int64               bias;
     int64               zero;
     int32               block_size;
     int32               n_blocks;
int64 int_bias;

The compression bias as a negative integer, e.g. if bias in the file header record is 100.0, then int_bias is −100 (this is the only value yet observed in practice).

int64 zero;

Always observed to be zero.

int32 block_size;

The number of bytes in each ZLIB compressed data block, except possibly the last, following decompression. Only 0x3ff000 has been observed so far.

int32 n_blocks;

The number of ZLIB compressed data blocks, always exactly

          (ztrailer_len - 24) / 24.

The fixed header is followed by n_blocks 24-byte ZLIB data block descriptors, each of which describes the compressed data block corresponding to its offset. Each block descriptor has the following format:

     int64               uncompressed_ofs;
     int64               compressed_ofs;
     int32               uncompressed_size;
     int32               compressed_size;
int64 uncompressed_ofs;

The offset, in bytes, that this block of data would have in a similar system file that uses compression format 1. This is zheader_ofs in the first block descriptor, and in each succeeding block descriptor it is the sum of the previous desciptor’s uncompressed_ofs and

          uncompressed_size.
int64 compressed_ofs;

The offset, in bytes, of the actual beginning of this compressed data block. This is zheader_ofs + 24 in the first block descriptor, and in each succeeding block descriptor it is the sum of the previ-

ous descriptor’s compressed_ofs and compressed_size. The final
block descriptor’s compressed_ofs and compressed_size sum to
ztrailer_ofs.
int32 uncompressed_size;

The number of bytes in this data block, after decompression. This is block_size in every data block except the last, which may be smaller.

int32 compressed_size;

The number of bytes in this data block, as stored compressed in this system file.

2 SPSS Viewer File Format

SPSS Viewer or .spv files, here called SPV files, are written by SPSS 16 and later to represent the contents of its output editor. This chapter documents the format, based on examination of a corpus of about 8,000 files from a variety of sources. This description is detailed enough to both read and write SPV files.

SPSS 15 and earlier versions instead use .spo files, which have a completely different output format based on the Microsoft Compound Document Format. This format is not documented here.

An SPV file is a Zip archive that can be read with zipinfo and unzip and similar programs. The final member in the Zip archive is the manifest, a file named META-INF/MANIFEST.MF. This structure makes SPV files resemble Java “JAR” files (and ODF files), but whereas a JAR manifest contains a sequence of colon-delimited key/value pairs, an SPV manifest contains the string ‘allowPivoting=true’, without a new-line. PSPP uses this string to identify an SPV file; it is invariant across the corpus.^{1}^{2}

The rest of the members in an SPV file’s Zip archive fall into two categories: structure and detail members. Structure member names take the form with

outputViewernumber.xml or outputViewernumber_heading.xml, where number is an

10-digit decimal number. Each of these members represents some kind of output item (a table, a heading, a block of text, etc.) or a group of them. The member whose output goes at the beginning of the document is numbered 0, the next member in the output is numbered 1, and so on.

Structure members contain XML. This XML is sometimes self-contained, but it often references detail members in the Zip archive, which are named as follows:

prefix_table.xml and prefix_tableData.bin
prefix_lightTableData.bin

The structure of a table plus its data. Older SPV files pair a prefix_table.xml file that describes the table’s structure with a binary prefix_tableData.bin file that gives its data. Newer SPV files (the majority of those in the corpus) instead include a single prefix_lightTableData.bin file that incorporates both into a single binary format.

prefix_warning.xml and prefix_warningData.bin
prefix_lightWarningData.bin

Same format used for tables, with a different name.

prefix_notes.xml and prefix_notesData.bin
prefix_lightNotesData.bin

Same format used for tables, with a different name.

prefix_chartData.bin and prefix_chart.xml

The structure of a chart plus its data. Charts do not have a “light” format.

1 SPV files always begin with the 7-byte sequence 50 4b 03 04 14 00 08, but this is not a useful magic

number because most Zip archives start the same way.

2 SPSS writes META-INF/MANIFEST.MF to every SPV file, but it does not read it or even require it to exist,

so using different contents, e.g. as ‘allowingPivot=false’ has no effect.

prefix_Imagegeneric.png
prefix_PastedObjectgeneric.png
prefix_imageData.bin

A PNG image referenced by an object element (in the first two cases) or an image element (in the final case). See Section 2.1.9 [SPV Structure object and image Elements], page 37.

prefix_pmml.scf
prefix_stats.scf
prefix_model.xml

Not yet investigated. The corpus contains few examples.

The prefix in the names of the detail members is typically an 11-digit decimal number that increases for each item, tending to skip values. Older SPV files use different naming conventions for detail members. Structure member refer to detail members by name, and so their exact names do not matter to readers as long as they are unique.

SPSS tolerates corrupted Zip archives that Zip reader libraries tend to reject. These can be fixed up with zip -FF.

2.1 Structure Member Format

A structure member lays out the high-level structure for a group of output items such as heading, tables, and charts. Structure members do not include the details of tables and charts but instead refer to them by their member names.

Structure members’ XML files claim conformance with a collection of XML Schemas. These schemas are distributed, under a nonfree license, with SPSS binaries. Fortunately, the schemas are not necessary to understand the structure members. The schemas can even be deceptive because they document elements and attributes that are not in the corpus and do not document elements and attributes that are commonly found in the corpus.

Structure members use a different XML namespace for each schema, but these namespaces are not entirely consistent. In some SPV files, for example, the viewer-tree schema is associated with namespace ‘http://xml.spss.com/spss/viewer-tree’ and in others

with ‘http://xml.spss.com/spss/viewer/viewer-tree’ (note the additional viewer/).

Under either name, the schema URIs are not resolvable to obtain the schemas themselves.

One may ignore all of the above in interpreting a structure member. The actual XML has a simple and straightforward form that does not require a reader to take schemas or namespaces into account. A structure member’s root is heading element, which contains heading or container elements (or a mix), forming a tree. In turn, container holds a label and one more child, usually text or table.

The following sections document the elements found in structure members in a contextfree grammar-like fashion. Consider the following example, which specifies the attributes and content for the container element:

     container
        :visibility=(visible | hidden)
        :page-break-before=(always)?
        :text-align=(left | center)?
        :width=dimension

     => label (table | container_text | graph | model | object | image | tree)

Each attribute specification begins with ‘:’ followed by the attribute’s name. If the attribute’s value has an easily specified form, then ‘=’ and its description follows the name. Finally, if the attribute is optional, the specification ends with ‘?’. The following value specifications are defined:

(a | b | ...)

One of the listed literal strings. If only one string is listed, it is the only acceptable value. If OTHER is listed, then any string not explicitly listed is also accepted.

bool  Either true or false.
dimension

A floating-point number followed by a unit, e.g. 10pt. Units in the corpus include in (inch), pt (points, 72/inch), px (“device-independent pixels”, 96/inch), and cm. If the unit is omitted then points should be assumed. The number and unit may be separated by white space. The corpus also includes localized names for units. A reader must understand these to properly interpret the dimension:

          inch  , pol., cala, cali

point

centimeter

real A floating-point number.

int An integer.

color A color in one of the forms #rrggbb or rrggbb, or the string transparent, or one of the standard Web color names.

ref
ref element
ref(elem1 | elem2 | ...)

The name from the id attribute in some element. If one or more elements are named, the name must refer to one of those elements, otherwise any element is acceptable.

All elements have an optional id attribute. If present, its value must be unique. In practice many elements are assigned id attributes that are never referenced.

The content specification for an element supports the following syntax:

element An element.

a b a followed by b.

a | b | c One of a or b or c.

a? Zero or one instances of a.

a* Zero or more instances of a. b+ One or more instances of a.

(subexpression)

Grouping for a subexpression.

EMPTY No content.

TEXT Text and CDATA.

Element and attribute names are sometimes suffixed by another name in square brackets to distinguish different uses of the same name. For example, structure XML has two text elements, one inside container, the other inside pageParagraph. The former is defined as text[container_text] and referenced as container_text, the latter defined as

text[pageParagraph_text] and referenced as pageParagraph_text.

This language is used in the PSPP source code for parsing structure and

detail XML members.  Refer to src/output/spv/structure-xml.grammar and
src/output/spv/detail-xml.grammar for the full grammars.

The following example shows the contents of a typical structure member for a DESCRIP-TIVES procedure. A real structure member is not indented. This example also omits most attributes, all XML namespace information, and the CSS from the embedded HTML:

     <?xml version="1.0" encoding="utf-8"?>
     <heading>
       <label>Output</label>
       <heading commandName="Descriptives">
         <label>Descriptives</label>
         <container>
           <label>Title</label>
           <text commandName="Descriptives" type="title">
             <html lang="en">
     <![CDATA[<head><style type="text/css">...</style></head><BR>Descriptives]]>
             </html>
           </text>
         </container>
         <container visibility="hidden">
           <label>Notes</label>
           <table commandName="Descriptives" subType="Notes" type="note">
             <tableStructure>
               <dataPath>00000000001_lightNotesData.bin</dataPath>
             </tableStructure>
           </table>
         </container>
         <container>
           <label>Descriptive Statistics</label>
           <table commandName="Descriptives" subType="Descriptive Statistics"
                  type="table">
             <tableStructure>
               <dataPath>00000000002_lightTableData.bin</dataPath>
             </tableStructure>

           </table>
         </container>
       </heading>
     </heading>

2.1.1 The heading Element

     heading[root_heading]
        :creator-version?
        :creator?
        :creation-date-time?
        :lockReader=bool?
        :schemaLocation?
     => label pageSetup? (container | heading)*
     heading
        :creator-version?
        :commandName?
        :visibility[heading_visibility]=(collapsed)?
        :locale?
        :olang?
     => label (container | heading)*

A heading represents a tree of content that appears in an output viewer window. It contains a label text string that is shown in the outline view ordinarily followed by content containers or further nested (sub)-sections of output. Unlike heading elements in HTML and other common document formats, which precede the content that they head, heading contains the elements that appear below the heading.

The root of a structure member is a special heading. The direct children of the root heading elements in all structure members in an SPV file are siblings. That is, the root heading in all of the structure members conceptually represent the same node. The root heading’s label is ignored (see see Section 2.1.2 [SPV Structure label Element], page 33). The root heading in the first structure member in the Zip file may contain a pageSetup element.

The schema implies that any heading may contain a sequence of any number of heading and container elements. This does not work for the root heading in practice, which must actually contain exactly one container or heading child element. Furthermore, if the root heading’s child is a heading, then the structure member’s name must end in _heading.xml; if it is a container child, then it must not.

The following attributes have been observed on both document root and nested heading elements.

[Attribute] creator-version

The version of the software that created this SPV file. A string of the form xxyyzzww represents software version xx.yy.zz.ww, e.g. 21000001 is version 21.0.0.1. Trailing pairs of zeros are sometimes omitted, so that 21, 210000, and 21000000 are all version 21.0.0.0 (and the corpus contains all three of those forms).

The following attributes have been observed on document root heading elements only:

[Attribute] creator

The directory in the file system of the software that created this SPV file.

[Attribute] creation-date-time

The date and time at which the SPV file was written, in a locale-specific format,

e.g. Friday, May 16, 2014 6:47:37 PM PDT or lunedì 17 marzo 2014 3.15.48 CET
or even Friday, December 5, 2014 5:00:19 o'clock PM EST.

[Attribute] lockReader

Whether a reader should be allowed to edit the output. The possible values are true and false. The value false is by far the most common.

[Attribute] schemaLocation

This is actually an XML Namespace attribute. A reader may ignore it.

The following attributes have been observed only on nested heading elements:

[Attribute] commandName

A locale-invariant identifier for the command that produced the output, e.g.

Frequencies, T-Test, Non Par Corr.

[Attribute] visibility

If this attribute is absent, the heading’s content is expanded in the outline view. If it is set to collapsed, it is collapsed. (This attribute is never present in a root heading because the root node is always expanded when a file is loaded, even though the UI can be used to collapse it interactively.)

[Attribute] locale

The locale used for output, in Windows format, which is similar to the format used in Unix with the underscore replaced by a hyphen, e.g. en-US, en-GB, el-GR, sr-Cryl-

RS.

[Attribute] olang

The output language, e.g. en, it, es, de, pt-BR.

2.1.2 The label Element

label => TEXT

Every heading and container holds a label as its first child. The label text is what appears in the outline pane of the GUI’s viewer window. PSPP also puts it into the outline of PDF output. The label text doesn’t appear in the output itself.

The text in label describes what it labels, often by naming the statistical procedure that was executed, e.g. “Frequencies” or “T-Test”. Labels are often very generic, especially within a container, e.g. “Title” or “Warnings” or “Notes”. Label text is localized according to the output language, e.g. in Italian a frequency table procedure is labeled “Frequenze”.

The user can edit labels to be anything they want. The corpus contains a few examples of empty labels, ones that contain no text, probably as a result of user editing.

The root heading in an SPV file has a label, like every heading. It normally contains “Output” but its content is disregarded anyway. The user cannot edit it.

2.1.3 The container Element

container
   :visibility=(visible | hidden)
   :page-break-before=(always)?
   :text-align=(left | center)?
   :width=dimension
=> label (table | container_text | graph | model | object | image | tree)

A container serves to contain and label a table, text, or other kind of item. This element has the following attributes.

[Attribute] visibility

Whether the container’s content is displayed. “Notes” tables are often hidden; other data is usually visible.

[Attribute] text-align

Alignment of text within the container. Observed with nested table and text elements.

[Attribute] width

The width of the container, e.g. 1097px.

All of the elements that nest inside container (except the label) have the following optional attribute.

[Attribute] commandName

As on the heading element. The corpus contains one example of where commandName is present but set to the empty string.

2.1.4 The text Element (Inside container)

text[container_text]
  :type[text_type]=(title | log | text | page-title)
  :commandName?
  :creator-version?
=> html

This text element is nested inside a container. There is a different text element that is nested inside a pageParagraph.

This element has the following attributes.

[Attribute] commandName

See Section 2.1.3 [SPV Structure container Element], page 34. For output not specific to a command, this is simply log.

[Attribute] type

The semantics of the text.

[Attribute] creator-version

As on the heading element.

2.1.5 The html Element

     html :lang=(en) => TEXT

The element contains an HTML document as text (or, in practice, as CDATA). In some cases, the document starts with <html> and ends with </html>; in others the html element is implied. Generally the HTML includes a head element with a CSS stylesheet. The HTML body often begins with <BR>.

The HTML document uses only the following elements:

html Sometimes, the document is enclosed with <html>…</html>.

br The HTML body often begins with <BR> and may contain it as well.

b
i

u Styling.

font The attributes face, color, and size are observed. The value of color takes one of the forms #rrggbb or rgb (r, g, b). The value of size is a number between 1 and 7, inclusive.

The CSS in the corpus is simple. To understand it, a parser only needs to be able to skip white space, <!--, and -->, and parse style only for p elements. Only the following properties matter:

color In the form rrggbb, e.g. 000000, with no leading ‘#’.

font-weight
          Either bold or normal.
font-style
          Either italic or normal.
text-decoration
          Either underline or normal.
font-family

A font name, commonly Monospaced or SansSerif.

font-size

Values claim to be in points, e.g. 14pt, but the values are actually in “deviceindependent pixels” (px), at 96/inch.

This element has the following attributes.

[Attribute] lang

This always contains en in the corpus.

2.1.6 The table Element

     table
        :VDPId?
        :ViZmlSource?
        :activePageId=int?
        :commandName

:creator-version?
:displayFiltering=bool?
:maxNumCells=int?
:orphanTolerance=int?
:rowBreakNumber=int?
:subType
:tableId
:tableLookId?
:type[table_type]=(table | note | warning)
=> tableProperties? tableStructure
tableStructure => path? dataPath csvPath?

This element has the following attributes.

[Attribute] commandName

See Section 2.1.3 [SPV Structure container Element], page 34.

[Attribute] type

One of table, note, or warning.

[Attribute] subType

The locale-invariant command ID for the particular kind of output that this table represents in the procedure. This can be the same as commandName e.g. Frequencies,

or different, e.g. Case Processing Summary. Generic subtypes Notes and Warnings

are often used.

[Attribute] tableId

A number that uniquely identifies the table within the SPV file, typically a large negative number such as -4147135649387905023.

[Attribute] creator-version

As on the heading element. In the corpus, this is only present for version 21 and up and always includes all 8 digits.

See Section 2.4.14 [SPV Detail Legacy Properties], page 81, for details on the

tableProperties element.

2.1.7 The graph Element

graph
:VDPId?
:ViZmlSource?
:commandName?
:creator-version?
:dataMapId?
:dataMapURI?
:editor?
:refMapId?
:refMapURI?

:csvFileIds?
:csvFileNames?
=> dataPath? path csvPath?

This element represents a graph. The dataPath and path elements name the Zip members that give the details of the graph. Normally, both elements are present; there is only one counterexample in the corpus.

csvPath only appears in one SPV file in the corpus, for two graphs. In these two cases, dataPath, path, and csvPath all appear. These csvPath name Zip members with names of the form number_csv.bin, where number is a many-digit number and the same as the csvFileIds. The named Zip members are CSV text files (despite the .bin extension). The CSV files are encoded in UTF-8 and begin with a U+FEFF byte-order marker.

2.1.8 The model Element

model
:PMMLContainerId?
:PMMLId
:StatXMLContainerId
:VDPId
:auxiliaryViewName
:commandName
:creator-version
:mainViewName
=> ViZml? dataPath? path | pmmlContainerPath statsContainerPath
pmmlContainerPath => TEXT
statsContainerPath => TEXT
ViZml :viewName? => TEXT

This element represents a model. The dataPath and path elements name the Zip members that give the details of the model. Normally, both elements are present; there is only one counterexample in the corpus.

The details are unexplored. The ViZml element contains base-64 encoded text, that decodes to a binary format with some embedded text strings, and path names an Zip member that contains XML. Alternatively, pmmlContainerPath and statsContainerPath name Zip members with .scf extension.

2.1.9 The object and image Elements

object
:commandName?
:type[object_type]=(unknown)?
:uri
=> EMPTY
image
:commandName?

        :VDPId
     => dataPath

These two elements represent an image in PNG format. They are equivalent and the corpus contains examples of both. The only difference is the syntax: for object, the uri attribute names the Zip member that contains a PNG file; for image, the text of the inner dataPath element names the Zip member.

PSPP writes object in output but there is no strong reason to choose this form.

The corpus only contains PNG image files.

2.1.10 The tree Element

     tree
        :commandName
        :creator-version
        :name
        :type
     => dataPath path

This element represents a tree. The dataPath and path elements name the Zip members that give the details of the tree. The details are unexplored.

2.1.11 Path Elements

     dataPath => TEXT
     path => TEXT
     csvPath => TEXT

These element contain the name of the Zip members that hold details for a container. For tables:

• When a “light” format is used, only dataPath is present, and it names a .bin member

    of the Zip file that has light in its name, e.g. 0000000001437_lightTableData.bin

(see Section 2.2 [SPV Light Detail Member Format], page 40).

• When the legacy format is used, both are present. In this case, dataPath names a Zip member with a legacy binary format that contains relevant data (see Section 2.3 [SPV Legacy Detail Member Binary Format], page 56), and path names a Zip member that uses an XML format (see Section 2.4 [SPV Legacy Detail Member XML Format], page 58).

Graphs normally follow the legacy approach described above. The corpus contains one example of a graph with path but not dataPath. The reason is unexplored.

Models use path but not dataPath. See Section 2.1.7 [SPV Structure graph Element], page 36, for more information.

These elements have no attributes.

2.1.12 The pageSetup Element

     pageSetup
        :initial-page-number=int?

   :chart-size=(as-is | full-height | half-height | quarter-height | OTHER)?
   :margin-left=dimension?
   :margin-right=dimension?
   :margin-top=dimension?
   :margin-bottom=dimension?
   :paper-height=dimension?
   :paper-width=dimension?
   :reference-orientation?
   :space-after=dimension?
=> pageHeader pageFooter
pageHeader => pageParagraph?
pageFooter => pageParagraph?
pageParagraph => pageParagraph_text

The pageSetup element has the following attributes.

[Attribute] initial-page-number

The page number to put on the first page of printed output. Usually 1.

[Attribute] chart-size

One of the listed, self-explanatory chart sizes, quarter-height, or a localization (!)

of one of these (e.g. dimensione attuale, Wie vorgegeben).
[Attribute] margin-left
[Attribute] margin-right

[Attribute] margin-top

[Attribute] margin-bottom

Margin sizes, e.g. 0.25in.

[Attribute] paper-height
[Attribute] paper-width

Paper sizes.

[Attribute] reference-orientation

Indicates the orientation of the output page. Either 0deg (portrait) or 90deg (landscape),

[Attribute] space-after

The amount of space between printed objects, typically 12pt.

2.1.13 The text Element (Inside pageParagraph)

text[pageParagraph_text] :type=(title | text) => TEXT

This text element is nested inside a pageParagraph. There is a different text element that is nested inside a container.

The element is either empty, or contains CDATA that holds almost-XHTML text: in the corpus, either an html or p element. It is almost-XHTML because the html element

designates the default namespace as ‘http://xml.spss.com/spss/viewer/viewer-tree’

instead of an XHTML namespace, and because the CDATA can contain substitution variables. The following variables are supported:

&[Date]

&[Time] The current date or time in the preferred format for the locale.

&[Head1]
&[Head2]
&[Head3]

&[Head4] First-, second-, third-, or fourth-level heading.

&[PageTitle]

The page title.

&[Filename]

Name of the output file.

&[Page] The page number.

&[Page] for the page number and &[PageTitle] for the page title.

Typical contents (indented for clarity):

     <html xmlns="http://xml.spss.com/spss/viewer/viewer-tree">
         <head></head>
         <body>
             <p style="text-align:right; margin-top: 0">Page &[Page]</p>
         </body>
     </html>

This element has the following attributes.

[Attribute] type

Always text.

2.2 Light Detail Member Format

This section describes the format of “light” detail .bin members. These members have a binary format which we describe here in terms of a context-free grammar using the following conventions:

NonTerminal ⇒ … Nonterminals have CamelCaps names, and ⇒ indicates a production. The right-hand side of a production is often broken across multiple lines. Break points are chosen for aesthetics only and have no semantic significance.

00, 01, …, ff.

A bytes with a fixed value, written as a pair of hexadecimal digits.

i0, i1, …, i9, i10, i11, … ib0, ib1, …, ib9, ib10, ib11, …

A 32-bit integer in little-endian or big-endian byte order, respectively, with a fixed value, written in decimal. Prefixed by ‘i’ for little-endian or ‘ib’ for big-endian. byte A byte.

bool A byte with value 0 or 1.

int16 be16 A 16-bit unsigned integer in little-endian or big-endian byte order, respectively.

int32 be32 A 32-bit unsigned integer in little-endian or big-endian byte order, respectively.

int64 be64 A 64-bit unsigned integer in little-endian or big-endian byte order, respectively.

double A 64-bit IEEE floating-point number.

float A 32-bit IEEE floating-point number.

string bestring A 32-bit unsigned integer, in little-endian or big-endian byte order, respectively,

followed by the specified number of bytes of character data. (The encoding is indicated by the Formats nonterminal.)

x? x is optional, e.g. 00? is an optional zero byte.

x*n x is repeated n times, e.g. byte*10 for ten arbitrary bytes.

x[name] Gives x the specified name. Names are used in textual explanations. They are also used, also bracketed, to indicate counts, e.g. int32[n] byte*[n] for a 32-bit integer followed by the specified number of arbitrary bytes.

a | b Either a or b.

(x) Parentheses are used for grouping to make precedence clear, especially in the presence of |, e.g. in 00 (01 | 02 | 03) 00.

count(x) becount(x)

A 32-bit unsigned integer, in little-endian or big-endian byte order, respectively, that indicates the number of bytes in x, followed by x itself.

v1(x) In a version 1 .bin member, x; in version 3, nothing. (The .bin header indicates the version.)

v3(x) In a version 3 .bin member, x; in version 1, nothing.

PSPP uses this grammar to parse light detail members. See src/output/spv/light-binary.grammar in the PSPP source tree for the full grammar.

Little-endian byte order is far more common in this format, but a few pieces of the format use big-endian byte order.

Light detail members express linear units in two ways: points (pt), at 72/inch, and “device-independent pixels” (px), at 96/inch. To convert from pt to px, multiply by 1.33 and round up. To convert from px to pt, divide by 1.33 and round down.

A “light” detail member .bin consists of a number of sections concatenated together, terminated by an optional byte 01:

     Table =>

Header Titles Footnotes
Areas Borders PrintSettings TableSettings Formats
Dimensions Axes Cells
01?

The following sections go into more detail.

2.2.1 Header

An SPV light member begins with a 39-byte header:

Header =>
01 00
(i1 | i3)[version]
bool[x0]
bool[x1]
bool[rotate-inner-column-labels]
bool[rotate-outer-row-labels]
bool[x2]
int32[x3]
int32[min-col-heading-width] int32[max-col-heading-width]
int32[min-row-heading-width] int32[max-row-heading-width]
int64[table-id]

version is a version number that affects the interpretation of some of the other data in the member. We will refer to “version 1” and “version 3” later on and use v1(…) and v3(…) for version-specific formatting (as described previously).

If rotate-inner-column-labels is 1, then column labels closest to the data are rotated 90˚ counterclockwise; otherwise, they are shown in the normal way.

If rotate-outer-row-labels is 1, then row labels farthest from the data are rotated 90˚ counterclockwise; otherwise, they are shown in the normal way.

min-col-heading-width, max-col-heading-width, min-row-heading-width, and

max-row-heading-width are measurements in 1/96 inch units (called “device independent pixel” units in Windows) whose values influence column widths. For the purpose of interpreting these values, a table is divided into the three regions shown below:

+------------------+-------------------------------------------------+
|                  |                  column headings                |
|                  +-------------------------------------------------+
|      corner      |                                                 |
|       and        |                                                 |
|   row headings   |                      data                       |
|                  |                                                 |
|                  |                                                 |
+------------------+-------------------------------------------------+
min-col-heading-width and max-col-heading-width apply to the columns in the col-

umn headings region. min-col-heading-width is the minimum width that any of these columns will be given automatically. In addition, max-col-heading-width is the maximum width that a column will be assigned to accommodate a long label in the column headings cells. These columns will still be made wider to accommodate wide data values in the data region.

min-row-heading-width is the minimum width that a column in the corner and row headings region will be given automatically. max-row-heading-width is the maximum width that a column in this region will be assigned to accomodate a long label. This region doesn’t include data, so data values don’t affect column widths.

table-id is a binary version of the tableId attribute in the structure member that refers to the detail member. For example, if tableId is -4122591256483201023, then table-id would be 0xc6c99d183b300001.

The meaning of the other variable parts of the header is not known. A writer may safely use version 3, true for x0, false for x1, true for x2, and 0x15 for x3.

2.2.2 Titles

     Titles =>
         Value[title] 01?
         Value[subtype] 01? 31
         Value[user-title] 01?
         (31 Value[corner-text] | 58)
         (31 Value[caption] | 58)

The Titles follow the Header and specify the table’s title, caption, and corner text.

The user-title reflects any user editing of the title text or style. The title is the title originally generated by the procedure. Both of these are appropriate for presentation and localized to the user’s language. For example, for a frequency table, title and user-title normally name the variable and c is simply “Frequencies”.

subtype is the same as the subType attribute in the table structure XML element that referred to this member. See Section 2.1.6 [SPV Structure table Element], page 35, for details.

The corner-text, if present, is shown in the upper-left corner of the table, above the row headings and to the left of the column headings. It is usually absent. When row dimension labels are displayed in the corner (see show-row-labels-in-corner), corner text is hidden.

The caption, if present, is shown below the table. caption reflects user editing of the caption.

2.2.3 Footnotes

     Footnotes => int32[n-footnotes] Footnote*[n-footnotes]
     Footnote => Value[text] (58 | 31 Value[marker]) int32[show]

Each footnote has text and an optional custom marker (such as ‘*’).

The syntax for Value would allow footnotes (and their markers) to reference other footnotes, but in practice this doesn’t work.

show is a 32-bit signed integer. It is positive to show the footnote or negative to hide it. Its magnitude is often 1, and in other cases tends to be the number of references to the footnote. It is safe to write 1 to show a footnote and -1 to hide it.

2.2.4 Areas

     Areas => 00? Area*8
     Area =>
         byte[index] 31
         string[typeface] float[size] int32[style] bool[underline]
         int32[halign] int32[valign]
         string[fg-color] string[bg-color]
         bool[alternate] string[alt-fg-color] string[alt-bg-color]
         v3(int32[left-margin] int32[right-margin] int32[top-margin] int32[bottom-margin])

Each Area represents the style for a different area of the table, in the following order: title, caption, footer, corner, column labels, row labels, data, and layers.

index is the 1-based index of the Area, i.e. 1 for the first Area, through 8 for the final Area.

typeface is the string name of the font used in the area. In the corpus, this is SansSerif in over 99% of instances and Times New Roman in the rest.

size is the size of the font, in px (see Section 2.2 [SPV Light Detail Member Format], page 40). The most common size in the corpus is 12 px. Even though size has a floatingpoint type, in the corpus its values are always integers.

style is a bit mask. Bit 0 (with value 1) is set for bold, bit 1 (with value 2) is set for italic.

underline is 1 if the font is underlined, 0 otherwise.

halign specifies horizontal alignment: 0 for center, 2 for left, 4 for right, 61453 for decimal, 64173 for mixed. Mixed alignment varies according to type: string data is leftjustified, numbers and most other formats are right-justified.

valign specifies vertical alignment: 0 for center, 1 for top, 3 for bottom.

fg-color and bg-color are the foreground color and background color, respectively. In the corpus, these are always #000000 and #ffffff, respectively.

alternate is 1 if rows should alternate colors, 0 if all rows should be the same color. When alternate is 1, alt-fg-color and alt-bg-color specify the colors for the alternate rows; otherwise they are empty strings.

   left-margin, right-margin, top-margin, and bottom-margin are measured in px.

2.2.5 Borders

     Borders =>
         count(
             ib1[endian]
             be32[n-borders] Border*[n-borders]
             bool[show-grid-lines]
             00 00 00)
     Border =>
         be32[border-type]
         be32[stroke-type]
         be32[color]

The Borders reflect how borders between regions are drawn.

The fixed value of endian can be used to validate the endianness.

show-grid-lines is 1 to draw grid lines, otherwise 0.

Each Border describes one kind of border. n-borders seems to always be 19. Each border-type appears once (although in an unpredictable order) and correspond to the following borders:

0 Title.

1…4 Left, top, right, and bottom outer frame.

5…8 Left, top, right, and bottom inner frame.

9, 10 Left and top of data area.

11, 12 Horizontal and vertical dimension rows.

13, 14 Horizontal and vertical dimension columns.

15, 16 Horizontal and vertical category rows.

17, 18 Horizontal and vertical category columns.

stroke-type describes how a border is drawn, as one of:

0 No line.

1 Solid line.

2 Dashed line.

3 Thick line.

4 Thin line.

5 Double line.

color is an RGB color. Bits 24–31 are alpha, bits 16–23 are red, 8–15 are green, 0–7 are blue. An alpha of 255 indicates an opaque color, therefore opaque black is 0xff000000.

2.2.6 Print Settings

     PrintSettings =>
         count(
             ib1[endian]
             bool[all-layers]
             bool[paginate-layers]
             bool[fit-width]
             bool[fit-length]
             bool[top-continuation]
             bool[bottom-continuation]
             be32[n-orphan-lines]
             bestring[continuation-string])

The PrintSettings reflect settings for printing. The fixed value of endian can be used to validate the endianness.

all-layers is 1 to print all layers, 0 to print only the layer designated by current-layer in TableSettings (see Section 2.2.7 [SPV Light Member Table Settings], page 46).

paginate-layers is 1 to print each layer at the start of a new page, 0 otherwise. (This setting is honored only all-layers is 1, since otherwise only one layer is printed.)

fit-width and fit-length control whether the table is shrunk to fit within a page’s width or length, respectively.

n-orphan-lines is the minimum number of rows or columns to put in one part of a table that is broken across pages.

If top-continuation is 1, then continuation-string is printed at the top of a page

when a table is broken across pages for printing; similarly for bottom-continuation and the bottom of a page. Usually, continuation-string is empty.

2.2.7 Table Settings

TableSettings =>
count(
v3(
ib1[endian]
be32[x5]
be32[current-layer]
bool[omit-empty]
bool[show-row-labels-in-corner]
bool[show-alphabetic-markers]
bool[footnote-marker-superscripts]
byte[x6]
becount(
  Breakpoints[row-breaks] Breakpoints[column-breaks]
  Keeps[row-keeps] Keeps[column-keeps]
  PointKeeps[row-point-keeps] PointKeeps[column-point-keeps]
)
bestring[notes]
bestring[table-look]
)...)
Breakpoints => be32[n-breaks] be32*[n-breaks]
Keeps => be32[n-keeps] Keep*[n-keeps]
Keep => be32[offset] be32[n]
PointKeeps => be32[n-point-keeps] PointKeep*[n-point-keeps]
PointKeep => be32[offset] be32 be32

The TableSettings reflect display settings. The fixed value of endian can be used to validate the endianness.

current-layer is the displayed layer. Suppose there are d layers, numbered 1 through d in the order given in the Dimensions (see Section 2.2.9 [SPV Light Member Dimensions], page 50), and that the displayed value of dimension i is d_{i}, 0 ≤ x_{i} < n_{i}, where n_{i} is the number of categories in dimension i. Then current-layer is calculated by the following algorithm:

     let current-layer = 0

for each i from d downto 1:

        current-layer = (n_{i} × current-layer) + x_{i}

If omit-empty is 1, empty rows or columns (ones with nothing in any cell) are hidden; otherwise, they are shown.

If show-row-labels-in-corner is 1, then row labels are shown in the upper left corner; otherwise, they are shown nested.

If show-alphabetic-markers is 1, markers are shown as letters (e.g. ‘a’, ‘b’, ‘c’, …); otherwise, they are shown as numbers starting from 1.

When footnote-marker-superscripts is 1, footnote markers are shown as superscripts, otherwise as subscripts.

The Breakpoints are rows or columns after which there is a page break; for example, a row break of 1 requests a page break after the second row. Usually no breakpoints are specified, indicating that page breaks should be selected automatically.

The Keeps are ranges of rows or columns to be kept together without a page break; for example, a row Keep with offset 1 and n 10 requests that the 10 rows starting with the second row be kept together. Usually no Keeps are specified.

The PointKeeps seem to be generated automatically based on user-specified Keeps. They seems to indicate a conversion from rows or columns to pixel or point offsets.

notes is a text string that contains user-specified notes. It is displayed when the user hovers the cursor over the table, like text in the title attribute in HTML. It is not printed. It is usually empty.

table-look is the name of a SPSS “TableLook” table style, such as “Default” or “Academic”; it is often empty.

TableSettings ends with an arbitrary number of null bytes. A writer may safely write 82 null bytes.

A writer may safely use 4 for x5 and 0 for x6.

2.2.8 Formats

     Formats =>
         int32[n-widths] int32*[n-widths]
         string[locale]
         int32[current-layer]
         bool[x7] bool[x8] bool[x9]
         Y0
         CustomCurrency
         count(
           v1(X0?)
           v3(count(X1 count(X2)) count(X3)))
     Y0 => int32[epoch] byte[decimal] byte[grouping]
     CustomCurrency => int32[n-ccs] string*[n-ccs]

If n-widths is nonzero, then the accompanying integers are column widths as manually adjusted by the user.

locale is a locale including an encoding, such as en_US.windows-1252 or it_IT.windows-1252. (locale is often duplicated in Y1, described below).

epoch is the year that starts the epoch. A 2-digit year is interpreted as belonging to the 100 years beginning at the epoch. The default epoch year is 69 years prior to the current year; thus, in 2017 this field by default contains 1948. In the corpus, epoch ranges from 1943 to 1948, plus some contain -1.

decimal is the decimal point character. The observed values are ‘.’ and ‘,’.

grouping is the grouping character. Usually, it is ‘,’ if decimal is ‘.’, and vice versa. Other observed values are ‘'’ (apostrophe), ‘ ’ (space), and zero (presumably indicating that digits should not be grouped).

n-ccs is observed as either 0 or 5. When it is 5, the following strings are CCA through CCE format strings. See Section “Custom Currency Formats” in PSPP. Most commonly these are all -,,, but other strings occur.

A writer may safely use false for x7, x8, and x9.

X0

X0 only appears, optionally, in version 1 members.

     X0 => byte*14 Y1 Y2
     Y1 =>
         string[command] string[command-local]
         string[language] string[charset] string[locale]
         bool[x10] bool[include-leading-zero] bool[x12] bool[x13]
         Y0
     Y2 => CustomCurrency byte[missing] bool[x17]

command describes the statistical procedure that generated the output, in English. It is not necessarily the literal syntax name of the procedure: for example, NPAR TESTS becomes “Nonparametric Tests.” command-local is the procedure’s name, translated into the output language; it is often empty and, when it is not, sometimes the same as command.

include-leading-zero is the LEADZERO setting for the table, where false is OFF (the default) and true is ON. See Section “SET LEADZERO” in PSPP.

missing is the character used to indicate that a cell contains a missing value. It is always observed as ‘.’.

A writer may safely use false for x10 and x17 and true for x12 and x13.

X1

X1 only appears in version 3 members.

     X1 =>
         bool[x14]
         byte[show-title]
         bool[x16]
         byte[lang]
         byte[show-variables]

         byte[show-values]
         int32[x18] int32[x19]
         00*17
         bool[x20]
         bool[show-caption]

lang may indicate the language in use. Some values seem to be 0: en, 1: de, 2: es, 3:

it, 5: ko, 6: pl, 8: zh-tw, 10: pt_BR, 11: fr.

show-variables determines how variables are displayed by default. A value of 1 means to display variable names, 2 to display variable labels when available, 3 to display both (name followed by label, separated by a space). The most common value is 0, which probably means to use a global default.

show-values is a similar setting for values. A value of 1 means to display the value, 2 to display the value label when available, 3 to display both. Again, the most common value is 0, which probably means to use a global default.

show-title is 1 to show the caption, 10 to hide it.

show-caption is true to show the caption, false to hide it.

A writer may safely use false for x14, false for x16, 0 for lang, -1 for x18 and x19, and false for x20.

X2

X2 only appears in version 3 members.

     X2 =>
         int32[n-row-heights] int32*[n-row-heights]
         int32[n-style-map] StyleMap*[n-style-map]
         int32[n-styles] StylePair*[n-styles]
         count((i0 i0)?)
     StyleMap => int64[cell-index] int16[style-index]

If present, n-row-heights and the accompanying integers are row heights as manually adjusted by the user.

The rest of X2 specifies styles for data cells. At first glance this is odd, because each data cell can have its own style embedded as part of the data, but in practice X2 specifies a style for a cell only if that cell is empty (and thus does not appear in the data at all). Each StyleMap specifies the index of a blank cell, calculated the same was as in the Cells (see Section 2.2.12 [SPV Light Member Cells], page 52), along with a 0-based index into the accompanying StylePair array.

A writer may safely omit the optional i0 i0 inside the count(...).

X3

X3 only appears in version 3 members.

     X3 =>
         01 00 byte[x21] 00 00 00
         Y1
         double[small] 01
         (string[dataset] string[datafile] i0 int32[date] i0)?

         Y2
         (int32[x22] i0 01?)?

small is a small real number. In the corpus, it overwhelmingly takes the value 0.0001, with zero occasionally seen. Nonzero numbers with format 40 (see Section 2.2.13 [SPV Light Member Value], page 53) whose magnitudes are smaller than displayed in scientific notation. (Thus, a small of zero prevents scientific notation from being chosen.)

dataset is the name of the dataset analyzed to produce the output, e.g. DataSet1, and datafile the name of the file it was read from, e.g. C:\Users\foo\bar.sav. The latter is sometimes the empty string.

date is a date, as seconds since the epoch, i.e. since January 1, 1970. Pivot tables within an SPV file often have dates a few minutes apart, so this is probably a creation date for the table rather than for the file.

Sometimes dataset, datafile, and date are present and other times they are absent. The reader can distinguish by assuming that they are present and then checking whether the presumptive dataset contains a null byte (a valid string never will).

x22 is usually 0 or 2000000.

A writer may safely use 4 for x21 and omit x22 and the other optional bytes at the end.

Encoding

Formats contains several indications of character encoding:

• locale in Formats itself.

• locale in Y1 (in version 1, Y1 is optionally nested inside X0; in version 3, Y1 is nested inside X3).

• charset in version 3, in Y1.

• lang in X1, in version 3.

charset, if present, is a good indication of character encoding, and in its absence the encoding suffix on locale in Formats will work.

locale in Y1 can be disregarded: it is normally the same as locale in Formats, and it is only present if charset is also.

lang is not helpful and should be ignored for character encoding purposes.

However, the corpus contains many examples of light members whose strings are encoded in UTF-8 despite declaring some other character set. Furthermore, the corpus contains several examples of light members in which some strings are encoded in UTF-8 (and contain multibyte characters) and other strings are encoded in another character set (and contain non-ASCII characters). PSPP treats any valid UTF-8 string as UTF-8 and only falls back to the declared encoding for strings that are not valid UTF-8.

The pspp-output program’s strings command can help analyze the encoding in an SPV light member. Use pspp-output --help-dev to see its usage.

2.2.9 Dimensions

A pivot table presents multidimensional data. A Dimension identifies the categories associated with each dimension.

     Dimensions => int32[n-dims] Dimension*[n-dims]

     Dimension =>
         Value[name] DimProperties
         int32[n-categories] Category*[n-categories]
     DimProperties =>
         byte[x1]
         byte[x2]
         int32[x3]
         bool[hide-dim-label]
         bool[hide-all-labels]
         01 int32[dim-index]

name is the name of the dimension, e.g. Variables, Statistics, or a variable name. The meanings of x1 and x3 are unknown. x1 is usually 0 but many other values have

been observed. A writer may safely use 0 for x1 and 2 for x3.

x2 is 0, 1, or 2. For a pivot table with L layer dimensions, R row dimensions, and C column dimensions, x2 is 2 for the first L dimensions, 0 for the next R dimensions, and 1 for the remaining C dimensions. This does not mean that the layer dimensions must be presented first, followed by the row dimensions, followed by the column dimensions—on the contrary, they are frequently in a different order—but x2 must follow this pattern to prevent the pivot table from being misinterpreted.

If hide-dim-label is 00, the pivot table displays a label for the dimension itself. Because usually the group and category labels are enough explanation, it is usually 01.

If hide-all-labels is 01, the pivot table omits all labels for the dimension, including group and category labels. It is usually 00. When hide-all-labels is 01, show-dim-label is ignored.

dim-index is usually the 0-based index of the dimension, e.g. 0 for the first dimension, 1 for the second, and so on. Sometimes it is -1. There is no visible difference. A writer may safely use the 0-based index.

2.2.10 Categories

Categories are arranged in a tree. Only the leaf nodes in the tree are really categories; the others just serve as grouping constructs.

     Category => Value[name] (Leaf | Group)
     Leaf => 00 00 00 i2 int32[leaf-index] i0
     Group =>
         bool[merge] 00 01 int32[x23]
         i-1 int32[n-subcategories] Category*[n-subcategories]

name is the name of the category (or group).

A Leaf represents a leaf category. The Leaf’s leaf-index is a nonnegative integer unique within the Dimension and less than n-categories in the Dimension. If the user does not sort or rearrange the categories, then leaf-index starts at 0 for the first Leaf in the dimension and increments by 1 with each successive Leaf. If the user does sorts or rearrange the categories, then the order of categories in the file reflects that change and leaf-index reflects the original order.

A dimension can have no leaf categories at all. A table that contains such a dimension necessarily has no data at all.

A Group is a group of nested categories. Usually a Group contains at least one Category, so that n-subcategories is positive, but Groups with zero subcategories have been observed.

If a Group’s merge is 00, the most common value, then the group is really a distinct group that should be represented as such in the visual representation and user interface. If merge is 01, the categories in this group should be shown and treated as if they were direct children of the group’s containing group (or if it has no parent group, then direct children of the dimension), and this group’s name is irrelevant and should not be displayed. (Merged groups can be nested!)

Writers need not use merged groups.

A Group’s x23 appears to be i2 when all of the categories within a group are leaf categories that directly represent data values for a variable (e.g. in a frequency table or crosstabulation, a group of values in a variable being tabulated) and i0 otherwise. A writer may safely write a constant 0 in this field.

2.2.11 Axes

After the dimensions come assignment of each dimension to one of the axes: layers, rows, and columns.

     Axes =>
         int32[n-layers] int32[n-rows] int32[n-columns]
         int32*[n-layers] int32*[n-rows] int32*[n-columns]

The values of n-layers, n-rows, and n-columns each specifies the number of dimensions displayed in layers, rows, and columns, respectively. Any of them may be zero. Their values sum to n-dimensions from Dimensions (see Section 2.2.9 [SPV Light Member Dimensions], page 50).

The following n-dimensions integers, in three groups, are a permutation of the 0-based dimension numbers. The first n-layers integers specify each of the dimensions represented by layers, the next n-rows integers specify the dimensions represented by rows, and the final n-columns integers specify the dimensions represented by columns. When there is more than one dimension of a given kind, the inner dimensions are given first. (For the layer axis, this means that the first dimension is at the bottom of the list and the last dimension is at the top when the current layer is displayed.)

2.2.12 Cells

The final part of an SPV light member contains the actual data.

     Cells => int32[n-cells] Cell*[n-cells]
     Cell => int64[index] v1(00?) Value

A Cell consists of an index and a Value. Suppose there are d dimensions, numbered 1 through d in the order given in the Dimensions previously, and that dimension i has n_{i} categories. Consider the cell at coordinates x_{i}, 1 ≤ i ≤ d, and note that 0 ≤ x_{i} < n_{i}. Then the index is calculated by the following algorithm:

let index = 0

for each i from 1 to d:

index = (n_{i} × index) + x_{i}

For example, suppose there are 3 dimensions with 3, 4, and 5 categories, respectively. The cell at coordinates (1, 2, 3) has index 5 × (4 × (3 × 0 + 1) + 2) + 3 = 33. Within a given dimension, the index is the leaf-index in a Leaf.

2.2.13 Value

Value is used throughout the SPV light member format. It boils down to a number or a string.

Value => 00? 00? 00? 00? RawValue
RawValue =>
01 ValueMod int32[format] double[x]
| 02 ValueMod int32[format] double[x]
string[var-name] string[value-label] byte[show]
| 03 string[local] ValueMod string[id] string[c] bool[fixed]
| 04 ValueMod int32[format] string[value-label] string[var-name]
byte[show] string[s]
| 05 ValueMod string[var-name] string[var-label] byte[show]
| 06 string[local] ValueMod string[id] string[c]
| ValueMod string[template] int32[n-args] Argument*[n-args]
Argument =>
i0 Value
| int32[x] i0 Value*[x]      /* x > 0 */

There are several possible encodings, which one can distinguish by the first nonzero byte in the encoding.

01 The numeric value x, intended to be presented to the user formatted according to format, which is about the same as the format described for system files (see [System File Output Formats], page 8). The exception is that format 40 is not MTIME but instead approximately a synonym for F format with a different rule for whether a value is shown in scientific notation: a value in format 40 is shown in scientific notation if and only if it is nonzero and its magnitude is less than small (see Section 2.2.8 [SPV Light Member Formats], page 47). Most commonly, format has width 40 (the maximum). An x with the maximum negative double value -DBL_MAX represents the systemmissing value SYSMIS. (HIGHEST and LOWEST have not been observed.) See Chapter 1 [System File Format], page 2, for more about these special values.

02 Similar to 01, with the additional information that x is a value of variable

var-name and has value label value-label. Both var-name and value-label

can be the empty string, the latter very commonly. show determines whether to show the numeric value or the value label. A value of 1 means to show the value, 2 to show the label, 3 to show both, and 0 means to use the default specified in show-values (see Section 2.2.8 [SPV Light Member Formats], page 47).

03 A text string, in two forms: c is in English, and sometimes abbreviated or obscure, and local is localized to the user’s locale. In an English-language locale, the two strings are often the same, and in the cases where they differ, local is more appropriate for a user interface, e.g. c of “Not a PxP table for MCN...” versus local of “Computed only for a PxP table, where P must be greater than 1.” c and local are always either both empty or both nonempty. id is a brief identifying string whose form seems to resemble a programming language identifier, e.g. cumulative_percent or factor_14. It is not unique. fixed is 00 for text taken from user input, such as syntax fragment, expressions, file names, data set names, and 01 for fixed text strings such as names of procedures or statistics. In the former case, id is always the empty string; in the latter case, id is still sometimes empty.

04 The string value s, intended to be presented to the user formatted according to format. The format for a string is not too interesting, and the corpus contains many clearly invalid formats like A16.39 or A255.127 or A134.1, so readers should probably entirely disregard the format. PSPP only checks format to distinguish AHEX format. s is a value of variable var-name and has value label value-label. var-name is never empty but value-label is commonly empty. show has the same meaning as in the encoding for 02.

05 Variable var-name with variable label var-label. In the corpus, var-name is rarely empty and var-label is often empty. show determines whether to show the variable name or the variable label. A value of 1 means to show the name, 2 to show the label, 3 to show both, and 0 means to use the default specified in show-variables (see Section 2.2.8 [SPV Light Member Formats], page 47).

06 Similar to type 03, with fixed assumed to be true.

otherwise When the first byte of a RawValue is not one of the above, the RawValue starts with a ValueMod, whose syntax is described in the next section. (A ValueMod always begins with byte 31 or 58.) This case is a template string, analogous to printf, followed by one or more Arguments, each of which has one or more values. The template string is copied directly into the output except for the following special syntax,

\%
\:
\[

\] Each of these expands to the character following ‘\\’, to escape characters that have special meaning in template strings. These are effective inside and outside the [...] syntax forms described below.

\n Expands to a new-line, inside or outside the [...] forms described below.

^i Expands to a formatted version of argument i, which must have only a single value. For example, ^1 expands to the first argument’s

          value.

[:a:]i Expands a for each of the values in i. a should contain one or more ^j conversions, which are drawn from the values for argument i in order. Some examples from the corpus:

[:^1:]1 All of the values for the first argument, concatenated.

[:^1\n:]1

Expands to the values for the first argument, each followed by a new-line.

[:^1 = ^2:]2

Expands to x = y where x is the second argument’s first value and y is its second value. (This would be used only if the argument has two values. If there were more values, the second and third values would be directly concatenated, which would look funny.)

[a:b:]i This extends the previous form so that the first values are expanded using a and later values are expanded using b. For an unknown reason, within a the ^j conversions are instead written as %j. Some examples from the corpus:

[%1:*^1:]1

Expands to all of the values for the first argument, separated by ‘*’.

[%1 = %2:, ^1 = ^2:]1

Given appropriate values for the first argument, ex-

pands to X = 1, Y = 2, Z = 3.
[%1:, ^1:]1

Given appropriate values, expands to 1, 2, 3.

The template string is localized to the user’s locale.

A writer may safely omit all of the optional 00 bytes at the beginning of a Value, except that it should write a single 00 byte before a templated Value.

2.2.14 ValueMod

A ValueMod can specify special modifications to a Value.

ValueMod =>
58
| 31
int32[n-refs] int16*[n-refs]
int32[n-subscripts] string*[n-subscripts]
v1(00 (i1 | i2) 00? 00? int32 00? 00?)
v3(count(TemplateString StylePair))
TemplateString => count((count((i0 (58 | 31 55))?) (58 | 31 string[id]))?)
StylePair =>
(31 FontStyle | 58)

         (31 CellStyle | 58)
     FontStyle =>
         bool[bold] bool[italic] bool[underline] bool[show]
         string[fg-color] string[bg-color]
         string[typeface] byte[size]
     CellStyle =>
         int32[halign] int32[valign] double[decimal-offset]
         int16[left-margin] int16[right-margin]
         int16[top-margin] int16[bottom-margin]

A ValueMod that begins with “31” specifies special modifications to a Value.

Each of the n-refs integers is a reference to a Footnote (see Section 2.2.3 [SPV Light Member Footnotes], page 43) by 0-based index. Footnote markers are shown appended to the main text of the Value, as superscripts or subscripts.

The subscripts, if present, are strings to append to the main text of the Value, as subscripts. Each subscript text is a brief indicator, e.g. ‘a’ or ‘b’, with its meaning indicated by the table caption. When multiple subscripts are present, they are displayed separated by commas.

The id inside the TemplateString, if present, is a template string for substitutions using the syntax explained previously. It appears to be an English-language version of the localized template string in the Value in which the Template is nested. A writer may safely omit the optional fixed data in TemplateString.

FontStyle and CellStyle, if present, change the style for this individual Value. In FontStyle, bold, italic, and underline control the particular style. show is ordinarily 1; if it is 0, then the cell data is not shown. fg-color and bg-color are strings in the format #rrggbb, e.g. #ff0000 for red or #ffffff for white. The empty string is occasionally observed also. The size is a font size in units of 1/128 inch.

In CellStyle, halign is 0 for center, 2 for left, 4 for right, 6 for decimal, 0xffffffad for mixed. For decimal alignment, decimal-offset is the decimal point’s offset from the right side of the cell, in pt (see Section 2.2 [SPV Light Detail Member Format], page 40). valign specifies vertical alignment: 0 for center, 1 for top, 3 for bottom. left-margin,

right-margin, top-margin, and bottom-margin are in pt.

2.3 Legacy Detail Member Binary Format

Whereas the light binary format represents everything about a given pivot table, the legacy binary format conceptually consists of a number of named sources, each of which consists of a number of named variables, each of which is a 1-dimensional array of numbers or strings or a mix. Thus, the legacy binary member format is quite simple.

This section uses the same context-free grammar notation as in the previous section, with the following additions:

vAF(x) In a version 0xaf legacy member, x; in other versions, nothing. (The legacy member header indicates the version; see below.)

vB0(x) In a version 0xb0 legacy member, x; in other versions, nothing.

A legacy detail member .bin has the following overall format:

     LegacyBinary =>
         00 byte[version] int16[n-sources] int32[member-size]
         Metadata*[n-sources]
         #Data*[n-sources]
         #Strings?

version is a version number that affects the interpretation of some of the other data in the member. Versions 0xaf and 0xb0 are known. We will refer to “version 0xaf” and “version 0xb0” members later on.

A legacy member consists of n-sources data sources, each of which has Metadata and Data.

member-size is the size of the legacy binary member, in bytes.

The Data and Strings above are commented out because the Metadata has some oddities that mean that the Data sometimes seems to start at an unexpected place. The following section goes into detail.

2.3.1 Metadata

     Metadata =>
         int32[n-values] int32[n-variables] int32[data-offset]
         vAF(byte*28[source-name])
         vB0(byte*64[source-name] int32[x])

A data source has n-variables variables, each with n-values data values.

source-name is a 28- or 64-byte string padded on the right with 0-bytes. The names that appear in the corpus are very generic: usually tableData for pivot table data or source0 for chart data.

A given Metadata’s data-offset is the offset, in bytes, from the beginning of the member to the start of the corresponding Data. This allows programs to skip to the beginning of the data for a particular source. In every case in the corpus, the Data follow the Metadata in the same order, but it is important to use data-offset instead of reading sequentially through the file because of the exception described below.

One SPV file in the corpus has legacy binary members with version 0xb0 but a 28byte source-name field (and only a single source). In practice, this means that the 64byte source-name used in version 0xb0 has a lot of 0-bytes in the middle followed by the variable-name of the following Data. As long as a reader treats the first 0-byte in the source-name as terminating the string, it can properly interpret these members.

The meaning of x in version 0xb0 is unknown.

2.3.2 Numeric Data

     Data => Variable*[n-variables]
     Variable => byte*288[variable-name] double*[n-values]

Data follow the Metadata in the legacy binary format, with sources in the same order (but readers should use the data-offset in Metadata records, rather than reading sequentially). Each Variable begins with a variable-name that generally indicates its role in the pivot table, e.g. “cell”, “cellFormat”, “dimension0categories”, “dimension0group0”, followed by the numeric data, one double per datum. A double with the maximum negative double -DBL_MAX represents the system-missing value SYSMIS.

2.3.3 String Data

     Strings => SourceMaps[maps] Labels
     SourceMaps => int32[n-maps] SourceMap*[n-maps]
     SourceMap => string[source-name] int32[n-variables] VariableMap*[n-variables]
     VariableMap => string[variable-name] int32[n-data] DatumMap*[n-data]
     DatumMap => int32[value-idx] int32[label-idx]
     Labels => int32[n-labels] Label*[n-labels]
     Label => int32[frequency] string[label]

Each variable may include a mix of numeric and string data values. If a legacy binary member contains any string data, Strings is present; otherwise, it ends just after the last Data element.

The string data overlays the numeric data. When a variable includes any string data, its Variable represents the string values with a SYSMIS or NaN placeholder. (Not all such values need be placeholders.)

Each SourceMap provides a mapping between SYSMIS or NaN values in source source-name and the string data that they represent. n-variables is the number of variables in the source that include string data. More precisely, it is the 1-based index of the last variable in the source that includes any string data; thus, it would be 4 if there are 5 variables and only the fourth one includes string data.

A VariableMap repeats its variable’s name, but variables are always present in the same order as the source, starting from the first variable, without skipping any even if they have no string values. Each VariableMap contains DatumMap nonterminals, each of which maps from a 0-based index within its variable’s data to a 0-based label index, e.g. pair value-idx = 2, label-idx = 3, means that the third data value (which must be SYSMIS or NaN) is to be replaced by the string of the fourth Label.

The labels themselves follow the pairs. The valuable part of each label is the string label. Each label also includes a frequency that reports the number of DatumMaps that reference it (although this is not useful).

2.4 Legacy Detail Member XML Format

The design of the detail XML format is not what one would end up with for describing pivot tables. This is because it is a special case of a much more general format (“visualization XML” or “VizML”) that can describe a wide range of visualizations. Most of this generality is overkill for tables, and so we end up with a funny subset of a general-purpose format.

An XML Schema for VizML is available, distributed with SPSS binaries, under a nonfree license. It contains documentation that is occasionally helpful.

This section describes the detail XML format using the same notation already used for the structure XML format (see Section 2.1 [SPV Structure Member Format], page 29). See src/output/spv/detail-xml.grammar in the PSPP source tree for the full grammar that it uses for parsing.

The important elements of the detail XML format are:

• Variables. See Section 2.4.2 [SPV Detail Variable Elements], page 61.

• Assignment of variables to axes. A variable can appear as columns, or rows, or layers. The faceting element and its sub-elements describe this assignment.

• Styles and other annotations.

This description is not detailed enough to write legacy tables. Instead, write tables in the light binary format.

2.4.1 The visualization Element

visualization
:creator
:date
:lang
:name
:style[style_ref]=ref style
:type
:version
:schemaLocation?
=> visualization_extension?
userSource
(sourceVariable | derivedVariable)+
categoricalDomain?
graph
labelFrame[lf1]*
container?
labelFrame[lf2]*
style+
layerController?
extension[visualization_extension]
:numRows=int?
:showGridline=bool?
:minWidthSet=(true)?
:maxWidthSet=(true)?
=> EMPTY
userSource :missing=(listwise | pairwise)? => EMPTY
categoricalDomain => variableReference simpleSort
simpleSort :method[sort_method]=(custom) => categoryOrder
container :style=ref style => container_extension? location+ labelFrame*

extension[container_extension] :combinedFootnotes=(true) => EMPTY
layerController
   :source=(tableData)
   :target=ref label?
=> EMPTY

The visualization element is the root of detail XML member. It has the following attributes:

[Attribute] creator

The version of the software that created this SPV file, as a string of the form xxyyzz, which represents software version xx.yy.zz, e.g. 160001 is version 16.0.1. The corpus includes major versions 16 through 19.

[Attribute] date

The date on the which the file was created, as a string of the form YYYY-MM-DD.

[Attribute] lang

The locale used for output, in Windows format, which is similar to the format used in Unix with the underscore replaced by a hyphen, e.g. en-US, en-GB, el-GR, sr-Cryl-

RS.

[Attribute] name

The title of the pivot table, localized to the output language.

[Attribute] style

The base style for the pivot table. In every example in the corpus, the style element has no attributes other than id.

[Attribute] type

A floating-point number. The meaning is unknown.

[Attribute] version

The visualization schema version number. In the corpus, the value is one of 2.4, 2.5, 2.7, and 2.8.

The userSource element has no visible effect. The extension element as a child of visualization has the following attributes.

[Attribute] numRows

An integer that presumably defines the number of rows in the displayed pivot table.

[Attribute] showGridline

Always set to false in the corpus.

[Attribute] minWidthSet
[Attribute] maxWidthSet

Always set to true in the corpus.

The extension element as a child of container has the following attribute

[Attribute] combinedFootnotes

Meaning unknown.

The categoricalDomain and simpleSort elements have no visible effect. The layerController element has no visible effect.

2.4.2 Variable Elements

A “variable” in detail XML is a 1-dimensional array of data. Each element of the array may, independently, have string or numeric content. All of the variables in a given detail XML member either have the same number of elements or have zero elements.

Two different elements define variables and their content:

sourceVariable

These variables’ data comes from the associated tableData.bin member.

derivedVariable

These variables are defined in terms of a mapping function from a source variable, or they are empty.

A variable named cell always exists. This variable holds the data displayed in the table. Variables in detail XML roughly correspond to the dimensions in a light detail member.

Each dimension has the following variables with stylized names, where n is a number for the dimension starting from 0:

dimensionncategories

The dimension’s leaf categories (see Section 2.2.10 [SPV Light Member Categories], page 51).

dimensionngroup0

Present only if the dimension’s categories are grouped, this variable holds the group labels for the categories. Grouping is inferred through adjacent identical labels. Categories that are not part of a group have empty-string data in this variable.

dimensionngroup1

Present only if the first-level groups are further grouped, this variable holds the labels for the second-level groups. There can be additional variables with further levels of grouping.

dimensionn

An empty variable.

Determining the data for a (non-empty) variable is a multi-step process:

1. Draw initial data from its source, for a sourceVariable, or from another named vari-

able, for a derivedVariable.

2. Apply mappings from valueMapEntry elements within the derivedVariable element, if any.

3. Apply mappings from relabel elements within a format or stringFormat element in

the sourceVariable or derivedVariable element, if any.

4. If the variable is a sourceVariable with a labelVariable attribute, and there were no mappings to apply in previous steps, then replace each element of the variable by the corresponding value in the label variable.

A single variable’s data can be modified in two of the steps, if both valueMapEntry and relabel are used. The following example from the corpus maps several integers to 2, then maps 2 in turn to the string “Input”:

<derivedVariable categorical="true" dependsOn="dimension0categories"
               id="dimension0group0map" value="map(dimension0group0)">
<stringFormat>
  <relabel from="2" to="Input"/>
  <relabel from="10" to="Missing Value Handling"/>
  <relabel from="14" to="Resources"/>
  <relabel from="0" to=""/>
  <relabel from="1" to=""/>
  <relabel from="13" to=""/>
</stringFormat>
<valueMapEntry from="2;3;5;6;7;8;9" to="2"/>
<valueMapEntry from="10;11" to="10"/>
<valueMapEntry from="14;15" to="14"/>
<valueMapEntry from="0" to="0"/>
<valueMapEntry from="1" to="1"/>
<valueMapEntry from="13" to="13"/>
</derivedVariable>

2.4.2.1 The sourceVariable Element

sourceVariable
 :id
 :categorical=(true)
 :source
 :domain=ref categoricalDomain?
 :sourceName
 :dependsOn=ref sourceVariable?
 :label?
 :labelVariable=ref sourceVariable?
=> variable_extension* (format | stringFormat)?

This element defines a variable whose data comes from the tableData.bin member that corresponds to this .xml.

This element has the following attributes.

[Attribute] id

An id is always present because this element exists to be referenced from other elements.

[Attribute] categorical

Always set to true.

[Attribute] source

Always set to tableData, the source-name in the corresponding tableData.bin

member (see Section 2.3.1 [SPV Legacy Member Metadata], page 57).

[Attribute] sourceName

The name of a variable within the source, corresponding to the variable-name in the tableData.bin member (see Section 2.3.2 [SPV Legacy Member Numeric Data], page 57).

[Attribute] label

The variable label, if any.

[Attribute] labelVariable

The variable-name of a variable whose string values correspond one-to-one with the values of this variable and are suitable for use as value labels.

[Attribute] dependsOn

This attribute doesn’t affect the display of a table.

2.4.2.2 The derivedVariable Element

derivedVariable
   :id
   :categorical=(true)
   :value
   :dependsOn=ref sourceVariable?
=> variable_extension* (format | stringFormat)? valueMapEntry*

Like sourceVariable, this element defines a variable whose values can be used elsewhere in the visualization. Instead of being read from a data source, the variable’s data are defined by a mathematical expression.

This element has the following attributes.

[Attribute] id

An id is always present because this element exists to be referenced from other elements.

[Attribute] categorical

Always set to true.

[Attribute] value

An expression that defines the variable’s value. In theory this could be an arbitrary expression in terms of constants, functions, and other variables, e.g. (var1 + var2)/2. In practice, the corpus contains only the following forms of expressions:

constant(0)
constant(variable)

All zeros. The reason why a variable is sometimes named is unknown. Sometimes the “variable name” has spaces in it.

map(variable)

Transforms the values in the named variable using the valueMapEntrys contained within the element.

[Attribute] dependsOn

This attribute doesn’t affect the display of a table.

2.4.2.3 The valueMapEntry Element

valueMapEntry :from :to => EMPTY

A valueMapEntry element defines a mapping from one or more values of a source expression to a target value. (In the corpus, the source expression is always just the name of a variable.) Each target value requires a separate valueMapEntry. If multiple source values map to the same target value, they can be combined or separate.

In the corpus, all of the source and target values are integers. valueMapEntry has the following attributes.

[Attribute] from

A source value, or multiple source values separated by semicolons, e.g. 0 or

13;14;15;16.

[Attribute] to

The target value, e.g. 0.

2.4.3 The extension Element

This is a general-purpose “extension” element. Readers that don’t understand a given extension should be able to safely ignore it. The attributes on this element, and their meanings, vary based on the context. Each known usage is described separately below. The current extensions use attributes exclusively, without any nested elements.

container Parent Element

extension[container_extension] :combinedFootnotes=(true) => EMPTY

With container as its parent element, extension has the following attributes.

[Attribute] combinedFootnotes

Always set to true in the corpus.

sourceVariable and derivedVariable Parent Element
extension[variable_extension] :from :helpId => EMPTY
With sourceVariable or derivedVariable as its parent element, extension has the

following attributes. A given parent element often contains several extension elements that specify the meaning of the source data’s variables or sources, e.g.

<extension from="0" helpId="corrected_model"/>
<extension from="3" helpId="error"/>
<extension from="4" helpId="total_9"/>
<extension from="5" helpId="corrected_total"/>

More commonly they are less helpful, e.g.

<extension from="0" helpId="notes"/>
<extension from="1" helpId="notes"/>
<extension from="2" helpId="notes"/>

<extension from="5" helpId="notes"/>
<extension from="6" helpId="notes"/>
<extension from="7" helpId="notes"/>
<extension from="8" helpId="notes"/>
<extension from="12" helpId="notes"/>
<extension from="13" helpId="no_help"/>
<extension from="14" helpId="notes"/>

[Attribute] from

An integer or a name like “dimension0”.

[Attribute] helpId

An identifier.

2.4.4 The graph Element

graph
   :cellStyle=ref style
   :style=ref style
=> location+ coordinates faceting facetLayout interval
coordinates => EMPTY

graph has the following attributes.

[Attribute] cellStyle [Attribute] style

Each of these is the id of a style element (see Section 2.4.12 [SPV Detail style Element], page 80). The former is the default style for individual cells, the latter for the entire table.

2.4.5 The location Element

location
   :part=(height | width | top | bottom | left | right)
   :method=(sizeToContent | attach | fixed | same)
   :min=dimension?
   :max=dimension?
   :target=ref (labelFrame | graph | container)?
   :value?
=> EMPTY

Each instance of this element specifies where some part of the table frame is located. All the examples in the corpus have four instances of this element, one for each of the parts height, width, left, and top. Some examples in the corpus add a fifth for part bottom, even though it is not clear how all of top, bottom, and height can be honored at the same time. In any case, location seems to have little importance in representing tables; a reader can safely ignore it.

[Attribute] part

The part of the table being located.

[Attribute] method

How the location is determined:

sizeToContent

Based on the natural size of the table. Observed only for parts height

          and width.

attach Based on the location specified in target. Observed only for parts top

          and bottom.

fixed Using the value in value. Observed only for parts top, bottom, and left.

same Same as the specified target. Observed only for part left.

[Attribute] min

Minimum size. Only observed with value 100pt. Only observed for part width.

[Dependent] target

Required when method is attach or same, not observed otherwise. This identifies an element to attach to. Observed with the ID of title, footnote, graph, and other elements.

[Dependent] value

Required when method is fixed, not observed otherwise. Observed values are 0%,

0px, 1px, and 3px on parts top and left, and 100% on part bottom.

2.4.6 The faceting Element

faceting => layer[layers1]* cross layer[layers2]*
cross => (unity | nest) (unity | nest)
unity => EMPTY
nest => variableReference[vars]+
variableReference :ref=ref (sourceVariable | derivedVariable) => EMPTY
layer
   :variable=ref (sourceVariable | derivedVariable)
   :value
   :visible=bool?
   :method[layer_method]=(nest)?
   :titleVisible=bool?
=> EMPTY

The faceting element describes the row, column, and layer structure of the table. Its cross child determines the row and column structure, and each layer child (if any) represents a layer. Layers may appear before or after cross.

The cross element describes the row and column structure of the table. It has exactly two children, the first of which describes the table’s columns and the second the table’s rows.

Each child is a nest element if the table has any dimensions along the axis in question, otherwise a unity element.

A nest element contains of one or more dimensions listed from innermost to outermost, each represented by variableReference child elements. Each variable in a dimension is listed in order. See Section 2.4.2 [SPV Detail Variable Elements], page 61, for information on the variables that comprise a dimension.

A nest can contain a single dimension, e.g.:

<nest>
  <variableReference ref="dimension0categories"/>
  <variableReference ref="dimension0group0"/>
  <variableReference ref="dimension0"/>
</nest>

A nest can contain multiple dimensions, e.g.:

<nest>
  <variableReference ref="dimension1categories"/>
  <variableReference ref="dimension1group0"/>
  <variableReference ref="dimension1"/>
  <variableReference ref="dimension0categories"/>
  <variableReference ref="dimension0"/>
</nest>

A nest may have no dimensions, in which case it still has one variableReference child, which references a derivedVariable whose value attribute is constant(0). In the corpus, such a derivedVariable has row or column, respectively, as its id. This is equivalent to using a unity element in place of nest.

A variableReference element refers to a variable through its ref attribute.

Each layer element represents a dimension, e.g.:

<layer value="0" variable="dimension0categories" visible="true"/>
<layer value="dimension0" variable="dimension0" visible="false"/>

layer has the following attributes.

[Attribute] variable

Refers to a sourceVariable or derivedVariable element.

[Attribute] value

The value to select. For a category variable, this is always 0; for a data variable, it is the same as the variable attribute.

[Attribute] visible

Whether the layer is visible. Generally, category layers are visible and data layers are not, but sometimes this attribute is omitted.

[Attribute] method

When present, this is always nest.

2.4.7 The facetLayout Element

facetLayout => tableLayout setCellProperties[scp1]*
               facetLevel+ setCellProperties[scp2]*
tableLayout
   :verticalTitlesInCorner=bool
   :style=ref style?
   :fitCells=(ticks both)?
=> EMPTY

The facetLayout element and its descendants control styling for the table. Its tableLayout child has the following attributes

[Attribute] verticalTitlesInCorner

If true, in the absence of corner text, row headings will be displayed in the corner.

[Attribute] style

Refers to a style element.

[Attribute] fitCells

Meaning unknown.

The facetLevel Element
facetLevel :level=int :gap=dimension? => axis
axis :style=ref style => label? majorTicks
majorTicks
   :labelAngle=int
   :length=dimension
   :style=ref style
   :tickFrameStyle=ref style
   :labelFrequency=int?
   :stagger=bool?
=> gridline?
gridline
   :style=ref style
   :zOrder=int
=> EMPTY
Each facetLevel describes a variableReference or layer, and a table has one

facetLevel element for each such element. For example, an SPV detail member that contains four variableReference elements and two layer elements will contain six

facetLevel elements.

In the corpus, facetLevel elements and the elements that they describe are always in the same order. The correspondence may also be observed in two other ways. First, one may use the level attribute, described below. Second, in the corpus, a facetLevel always has an id that is the same as the id of the element it describes with _facetLevel appended. One should not formally rely on this, of course, but it is usefully indicative.

[Attribute] level

A 1-based index into the variableReference and layer elements, e.g. a facetLayout with a level of 1 describes the first variableReference in the SPV detail member, and in a member with four variableReference elements, a facetLayout with a level of 5 describes the first layer in the member.

[Attribute] gap

Always observed as 0pt.

Each facetLevel contains an axis, which in turn may contain a label for the facetLevel (see Section 2.4.8 [SPV Detail label Element], page 69) and does contain a

majorTicks element.

[Attribute] labelAngle

Normally 0. The value -90 causes inner column or outer row labels to be rotated vertically.

[Attribute] style

[Attribute] tickFrameStyle

Each refers to a style element. style is the style of the tick labels, tickFrameStyle the style for the frames around the labels.

2.4.8 The label Element

label
   :style=ref style
   :textFrameStyle=ref style?
   :purpose=(title | subTitle | subSubTitle | layer | footnote)?
=> text+ | descriptionGroup
descriptionGroup
   :target=ref faceting
   :separator?
=> (description | text)+
description :name=(variable | value) => EMPTY
text
   :usesReference=int?
   :definesReference=int?
   :position=(subscript | superscript)?
   :style=ref style
=> TEXT

This element represents a label on some aspect of the table.

[Attribute] style

[Attribute] textFrameStyle

Each of these refers to a style element. style is the style of the label text, textFrameStyle the style for the frame around the label.

[Attribute] purpose

The kind of entity being labeled.

A descriptionGroup concatenates one or more elements to form a label. Each element can be a text element, which contains literal text, or a description element that substitutes a value or a variable name.

[Attribute] target

The id of an element being described. In the corpus, this is always faceting.

[Attribute] separator

A string to separate the description of multiple groups, if the target has more than one. In the corpus, this is always a new-line.

Typical contents for a descriptionGroup are a value by itself:

<description name="value"/>

or a variable and its value, separated by a colon:

<description name="variable"/><text>:</text><description name="value"/>

A description is like a macro that expands to some property of the target of its parent descriptionGroup. The name attribute specifies the property.

2.4.9 The setCellProperties Element
setCellProperties
   :applyToConverse=bool?
=> (setStyle | setFrameStyle | setFormat | setMetaData)* union[union_]?

The setCellProperties element sets style properties of cells or row or column labels. Interpreting setCellProperties requires answering two questions: which cells or labels

to style, and what styles to use.

Which Cells?

union => intersect+
intersect => where+ | intersectWhere | alternating | EMPTY
where
   :variable=ref (sourceVariable | derivedVariable)
   :include
=> EMPTY
intersectWhere
   :variable=ref (sourceVariable | derivedVariable)
   :variable2=ref (sourceVariable | derivedVariable)

=> EMPTY
alternating => EMPTY

When union is present with intersect children, each of those children specifies a group of cells that should be styled, and the total group is all those cells taken together. When union is absent, every cell is styled. One attribute on setCellProperties affects the choice of cells:

[Attribute] applyToConverse

If true, this inverts the meaning of the cell selection: the selected cells are the ones not designated. This is confusing, given the additional restrictions of union, but in the corpus applyToConverse is never present along with union.

An intersect specifies restrictions on the cells to be matched. Each where child specifies which values of a given variable to include. The attributes of intersect are:

[Attribute] variable

Refers to a variable, e.g. dimension0categories. Only “categories” variables make sense here, but other variables, e.g. dimension0group0map, are sometimes seen. The reader may ignore these.

[Attribute] include

A value, or multiple values separated by semicolons, e.g. 0 or 13;14;15;16.

PSPP ignores setCellProperties when intersectWhere is present.

What Styles?

setStyle
   :target=ref (labeling | graph | interval | majorTicks)
   :style=ref style
=> EMPTY
setMetaData :target=ref graph :key :value => EMPTY
setFormat
   :target=ref (majorTicks | labeling)
   :reset=bool?
=> format | numberFormat | stringFormat+ | dateTimeFormat | elapsedTimeFormat
setFrameStyle
   :style=ref style
   :target=ref majorTicks
=> EMPTY

The set* children of setCellProperties determine the styles to set.

When setCellProperties contains a setFormat whose target references a labeling

element, or if it contains a setStyle that references a labeling or interval element, the setCellProperties sets the style for table cells. The format from the setFormat, if present, replaces the cells’ format. The style from the setStyle that references labeling, if present, replaces the label’s font and cell styles, except that the background color is taken instead from the interval’s style, if present.

   When setCellProperties contains a setFormat whose target references a majorTicks

element, or if it contains a setStyle whose target references a majorTicks, or if it contains a setFrameStyle element, the setCellProperties sets the style for row or column labels. In this case, the setCellProperties always contains a single where element whose variable designates the variable whose labels are to be styled. The format from the setFormat, if present, replaces the labels’ format. The style from the setStyle that references majorTicks, if present, replaces the labels’ font and cell styles, except that the background color is taken instead from the setFrameStyle’s style, if present.

   When setCellProperties contains a setStyle whose target references a graph el-

ement, and one that references a labeling element, and the union element contains alternating, the setCellProperties sets the alternate foreground and background colors for the data area. The foreground color is taken from the style referenced by the setStyle that targets the graph, the background color from the setStyle for labeling.

A reader may ignore a setCellProperties that only contains setMetaData, as well as

setMetaData within other setCellProperties.

A reader may ignore a setCellProperties whose only set* child is a setStyle that targets the graph element.

The setStyle Element

     setStyle
        :target=ref (labeling | graph | interval | majorTicks)
        :style=ref style
     => EMPTY

This element associates a style with the target.

[Attribute] target

The id of an element whose style is to be set.

[Attribute] style

The id of a style element that identifies the style to set on the target.

2.4.10 The setFormat Element

     setFormat
        :target=ref (majorTicks | labeling)
        :reset=bool?
     => format | numberFormat | stringFormat+ | dateTimeFormat | elapsedTimeFormat

This element sets the format of the target, “format” in this case meaning the SPSS print format for a variable.

The details of this element vary depending on the schema version, as declared in the root visualization element’s version attribute (see Section 2.4.1 [SPV Detail visualization Element], page 59). A reader can interpret the content without knowing the schema version.

The setFormat element itself has the following attributes.

[Attribute] target

Refers to an element whose style is to be set.

[Attribute] reset

If this is true, this format replaces the target’s previous format. If it is false, the modifies the previous format.

2.4.10.1 The numberFormat Element

numberFormat
   :minimumIntegerDigits=int?
   :maximumFractionDigits=int?
   :minimumFractionDigits=int?
   :useGrouping=bool?
   :scientific=(onlyForSmall | whenNeeded | true | false)?
   :small=real?
   :prefix?
   :suffix?
=> affix*

Specifies a format for displaying a number. The available options are a superset of those available from PSPP print formats. PSPP chooses a print format type for a numberFormat as follows:

1. If scientific is true, uses E format.

2. If prefix is $, uses DOLLAR format.

3. If suffix is %, uses PCT format.

4. If useGrouping is true, uses COMMA format.

5. Otherwise, uses F format.

For translating to a print format, PSPP uses maximumFractionDigits as the number of decimals, unless that attribute is missing or out of the range [0,15], in which case it uses 2 decimals.

[Attribute] minimumIntegerDigits

Minimum number of digits to display before the decimal point. Always observed as

0.
[Attribute] maximumFractionDigits
[Attribute] minimumFractionDigits

Maximum or minimum, respectively, number of digits to display after the decimal point. The observed values of each attribute range from 0 to 9.

[Attribute] useGrouping

Whether to use the grouping character to group digits in large numbers.

[Attribute] scientific

This attribute controls when and whether the number is formatted in scientific notation. It takes the following values:

onlyForSmall

Use scientific notation only when the number’s magnitude is smaller than the value of the small attribute.

whenNeeded

Use scientific notation when the number will not otherwise fit in the available space.

true Always use scientific notation. Not observed in the corpus.

false Never use scientific notation. A number that won’t otherwise fit will be replaced by an error indication (see the errorCharacter attribute). Not observed in the corpus.

[Attribute] small

Only present when the scientific attribute is onlyForSmall, this is a numeric magnitude below which the number will be formatted in scientific notation. The values 0 and 0.0001 have been observed. The value 0 seems like a pathological choice, since no real number has a magnitude less than 0; perhaps in practice such a choice is equivalent to setting scientific to false.

[Attribute] prefix [Attribute] suffix

Specifies a prefix or a suffix to apply to the formatted number. Only suffix has been observed, with value ‘%’.

2.4.10.2 The stringFormat Element

stringFormat => relabel* affix*
relabel :from=real :to => EMPTY

The stringFormat element specifies how to display a string. By default, a string is displayed verbatim, but relabel can change it.

The relabel element appears as a child of stringFormat (and of format, when it is used to format strings). It specifies how to display a given value. It is used to implement value labels and to display the system-missing value in a human-readable way. It has the following attributes:

[Attribute] from

The value to map. In the corpus this is an integer or the system-missing value

-1.797693134862316E300.

[Attribute] to

The string to display in place of the value of from. In the corpus this is a wide variety of value labels; the system-missing value is mapped to ‘.’.

2.4.10.3 The dateTimeFormat Element

dateTimeFormat
   :baseFormat[dt_base_format]=(date | time | dateTime)
   :separatorChars?
   :mdyOrder=(dayMonthYear | monthDayYear | yearMonthDay)?
   :showYear=bool?
   :yearAbbreviation=bool?
   :showQuarter=bool?

:quarterPrefix?
:quarterSuffix?
:showMonth=bool?
:monthFormat=(long | short | number | paddedNumber)?
:showWeek=bool?
:weekPadding=bool?
:weekSuffix?
:showDayOfWeek=bool?
:dayOfWeekAbbreviation=bool?
:dayPadding=bool?
:dayOfMonthPadding=bool?
:hourPadding=bool?
:minutePadding=bool?
:secondPadding=bool?
:showDay=bool?
:showHour=bool?
:showMinute=bool?
:showSecond=bool?
:showMillis=bool?
:dayType=(month | year)?
:hourFormat=(AMPM | AS_24 | AS_12)?
=> affix*

This element appears only in schema version 2.5 and earlier (see Section 2.4.1 [SPV Detail visualization Element], page 59).

Data to be formatted in date formats is stored as strings in legacy data, in the format yyyy-mm-ddTHH:MM:SS.SSS and must be parsed and reformatted by the reader.

The following attribute is required.

[Attribute] baseFormat

Specifies whether a date and time are both to be displayed, or just one of them.

Many of the attributes’ meanings are obvious. The following seem to be worth documenting.

[Attribute] separatorChars

Exactly four characters. In order, these are used for: decimal point, grouping, date separator, time separator. Always ‘.,-:’.

[Attribute] mdyOrder

Within a date, the order of the days, months, and years. dayMonthYear is the only observed value, but one would expect that monthDayYear and yearMonthDay to be reasonable as well.

[Attribute] showYear

[Attribute] yearAbbreviation

Whether to include the year and, if so, whether the year should be shown abbreviated, that is, with only 2 digits. Each is true or false; only values of true and false, respectively, have been observed.

[Attribute] showMonth

[Attribute] monthFormat

Whether to include the month (true or false) and, if so, how to format it. monthFormat is one of the following:

long The full name of the month, e.g. in an English locale, September.

short The abbreviated name of the month, e.g. in an English locale, Sep.

number The number representing the month, e.g. 9 for September.

paddedNumber

A two-digit number representing the month, e.g. 09 for September.

Only values of true and short, respectively, have been observed.

[Attribute] dayType

This attribute is always month in the corpus, specifying that the day of the month is to be displayed; a value of year is supposed to indicate that the day of the year, where 1 is January 1, is to be displayed instead.

[Attribute] hourFormat

hourFormat, if present, is one of:

AMPM The time is displayed with an am or pm suffix, e.g. 10:15pm.

AS_24 The time is displayed in a 24-hour format, e.g. 22:15.

This is the only value observed in the corpus.

AS_12 The time is displayed in a 12-hour format, without distinguishing morning or evening, e.g. 10;15.

hourFormat is sometimes present for elapsedTime formats, which is confusing since a time duration does not have a concept of AM or PM. This might indicate a bug in the code that generated the XML in the corpus, or it might indicate that elapsedTime is sometimes used to format a time of day.

For a baseFormat of date, PSPP chooses a print format type based on the following rules:

1. If showQuarter is true: QYR.

2. Otherwise, if showWeek is true: WKYR.

3. Otherwise, if mdyOrder is dayMonthYear:
a. If monthFormat is number or paddedNumber: EDATE.

b. Otherwise: DATE.

4. Otherwise, if mdyOrder is yearMonthDay: SDATE.

5. Otherwise, ADATE.

For a baseFormat of dateTime, PSPP uses YMDHMS if mdyOrder is yearMonthDay and
DATETIME otherwise. For a baseFormat of time, PSPP uses DTIME if showDay is true,

otherwise TIME if showHour is true, otherwise MTIME.

For a baseFormat of date, the chosen width is the minimum for the format type, adding 2 if yearAbbreviation is false or omitted. For other base formats, the chosen width is the minimum for its type, plus 3 if showSecond is true, plus 4 more if showMillis is also true. Decimals are 0 by default, or 3 if showMillis is true.

2.4.10.4 The elapsedTimeFormat Element

elapsedTimeFormat
:baseFormat[dt_base_format]=(date | time | dateTime)
:dayPadding=bool?
:hourPadding=bool?
:minutePadding=bool?
:secondPadding=bool?
:showYear=bool?
:showDay=bool?
:showHour=bool?
:showMinute=bool?
:showSecond=bool?
:showMillis=bool?
=> affix*

This element specifies the way to display a time duration. Data to be formatted in elapsed time formats is stored as strings in legacy data, in the

format H:MM:SS.SSS, with additional hour digits as needed for long durations, and must be parsed and reformatted by the reader.

The following attribute is required.

[Attribute] baseFormat

Specifies whether a day and a time are both to be displayed, or just one of them.

The remaining attributes specify exactly how to display the elapsed time. For baseFormat of time, PSPP converts this element to print format type DTIME; other-

wise, if showHour is true, to TIME; otherwise, to MTIME. The chosen width is the minimum for the chosen type, adding 3 if showSecond is true, adding 4 more if showMillis is also true. Decimals are 0 by default, or 3 if showMillis is true.

2.4.10.5 The format Element

format
:baseFormat[f_base_format]=(date | time | dateTime | elapsedTime)?
:errorCharacter?
:separatorChars?
:mdyOrder=(dayMonthYear | monthDayYear | yearMonthDay)?
:showYear=bool?
:showQuarter=bool?
:quarterPrefix?
:quarterSuffix?
:yearAbbreviation=bool?
:showMonth=bool?
:monthFormat=(long | short | number | paddedNumber)?
:dayPadding=bool?
:dayOfMonthPadding=bool?
:showWeek=bool?
:weekPadding=bool?

:weekSuffix?
:showDayOfWeek=bool?
:dayOfWeekAbbreviation=bool?
:hourPadding=bool?
:minutePadding=bool?
:secondPadding=bool?
:showDay=bool?
:showHour=bool?
:showMinute=bool?
:showSecond=bool?
:showMillis=bool?
:dayType=(month | year)?
:hourFormat=(AMPM | AS_24 | AS_12)?
:minimumIntegerDigits=int?
:maximumFractionDigits=int?
:minimumFractionDigits=int?
:useGrouping=bool?
:scientific=(onlyForSmall | whenNeeded | true | false)?
:small=real?
:prefix?
:suffix?
:tryStringsAsNumbers=bool?
:negativesOutside=bool?
=> relabel* affix*

This element is the union of all of the more-specific format elements. It is interpreted in the same way as one of those format elements, using baseFormat to determine which kind of format to use.

There are a few attributes not present in the more specific formats:

[Attribute] tryStringsAsNumbers

When this is true, it is supposed to indicate that string values should be parsed as numbers and then displayed according to numeric formatting rules. However, in the corpus it is always false.

[Attribute] negativesOutside

If true, the negative sign should be shown before the prefix; if false, it should be shown after.

2.4.10.6 The affix Element

affix
:definesReference=int
:position=(subscript | superscript)
:suffix=bool
:value
=> EMPTY

This defines a suffix (or, theoretically, a prefix) for a formatted value. It is used to insert a reference to a footnote. It has the following attributes:

[Attribute] definesReference

This specifies the footnote number as a natural number: 1 for the first footnote, 2 for the second, and so on.

[Attribute] position

Position for the footnote label. Always superscript.

[Attribute] suffix

Whether the affix is a suffix (true) or a prefix (false). Always true.

[Attribute] value

The text of the suffix or prefix. Typically a letter, e.g. a for footnote 1, b for footnote 2, … The corpus contains other values: *, **, and a few that begin with at least one comma: ,b, ,c, ,,b, and ,,c.

2.4.11 The interval Element

interval :style=ref style => labeling footnotes?
labeling
   :style=ref style?
   :variable=ref (sourceVariable | derivedVariable)
=> (formatting | format | footnotes)*
formatting :variable=ref (sourceVariable | derivedVariable) => formatMapping*
formatMapping :from=int => format?
footnotes
   :superscript=bool?
   :variable=ref (sourceVariable | derivedVariable)
=> footnoteMapping*
footnoteMapping :definesReference=int :from=int :to => EMPTY

The interval element and its descendants determine the basic formatting and labeling for the table’s cells. These basic styles are overridden by more specific styles set using setCellProperties (see Section 2.4.9 [SPV Detail setCellProperties Element], page 70).

The style attribute of interval itself may be ignored.

The labeling element may have a single formatting child. If present, its variable attribute refers to a variable whose values are format specifiers as numbers, e.g. value 0x050802 for F8.2. However, the numbers are not actually interpreted that way. Instead, each number actually present in the variable’s data is mapped by a formatMapping child of formatting to a format that specifies how to display it.

The labeling element may also have a footnotes child element. The variable attribute of this element refers to a variable whose values are comma-delimited strings that list the 1-based indexes of footnote references. (Cells without any footnote references are numeric 0 instead of strings.)

Each footnoteMapping child of the footnotes element defines the footnote marker to be its to attribute text for the footnote whose 1-based index is given in its definesReference attribute.

2.4.12 The style Element

style
:color=color?
:color2=color?
:labelAngle=real?
:border-bottom=(solid | thick | thin | double | none)?
:border-top=(solid | thick | thin | double | none)?
:border-left=(solid | thick | thin | double | none)?
:border-right=(solid | thick | thin | double | none)?
:border-bottom-color?
:border-top-color?
:border-left-color?
:border-right-color?
:font-family?
:font-size?
:font-weight=(regular | bold)?
:font-style=(regular | italic)?
:font-underline=(none | underline)?
:margin-bottom=dimension?
:margin-left=dimension?
:margin-right=dimension?
:margin-top=dimension?
:textAlignment=(left | right | center | decimal | mixed)?
:labelLocationHorizontal=(positive | negative | center)?
:labelLocationVertical=(positive | negative | center)?
:decimal-offset=dimension?
:size?
:width?
:visible=bool?
=> EMPTY

A style element has an effect only when it is referenced by another element to set some aspect of the table’s style. Most of the attributes are self-explanatory. The rest are described below.

[Attribute] color

In some cases, the text color; in others, the background color.

[Attribute] color2

Not used.

[Attribute] labelAngle

Normally 0. The value -90 causes inner column or outer row labels to be rotated vertically.

[Attribute] labelLocationHorizontal

Not used.

[Attribute] labelLocationVertical

The value positive corresponds to vertically aligning text to the top of a cell,

negative to the bottom, center to the middle.

2.4.13 The labelFrame Element

     labelFrame :style=ref style => location+ label? paragraph?
     paragraph :hangingIndent=dimension? => EMPTY

A labelFrame element specifies content and style for some aspect of a table. Only labelFrame elements that have a label child are important. The purpose attribute in the label determines what the labelFrame affects:

title The table’s title and its style.

subTitle The table’s caption and its style.

footnote The table’s footnotes and the style for the footer area.

layer The style for the layer area.

subSubTitle

Ignored.

The style attribute references the style to use for the area.

The label, if present, specifies the text to put into the title or caption or footnotes. For footnotes, the label has two text children for every footnote, each of which has a usesReference attribute identifying the 1-based index of a footnote. The first, third, fifth, … text child specifies the content for a footnote; the second, fourth, sixth, … child specifies the marker. Content tends to end in a new-line, which the reader may wish to trim; similarly, markers tend to end in ‘.’.

The paragraph, if present, may be ignored, since it is always empty.

2.4.14 Legacy Properties

The detail XML format has features for styling most of the aspects of a table. It also inherits defaults for many aspects from structure XML, which has the following tableProperties element:

     tableProperties
        :name?
     => generalProperties footnoteProperties cellFormatProperties borderProperties printingProperties
     generalProperties
        :hideEmptyRows=bool?
        :maximumColumnWidth=dimension?
        :maximumRowWidth=dimension?
        :minimumColumnWidth=dimension?
        :minimumRowWidth=dimension?

:rowDimensionLabels=(inCorner | nested)?
=> EMPTY
footnoteProperties
:markerPosition=(superscript | subscript)?
:numberFormat=(alphabetic | numeric)?
=> EMPTY
cellFormatProperties => cell_style+
any[cell_style]
:alternatingColor=color?
:alternatingTextColor=color?
=> style
style
:color=color?
:color2=color?
:font-family?
:font-size?
:font-style=(regular | italic)?
:font-weight=(regular | bold)?
:font-underline=(none | underline)?
:labelLocationVertical=(positive | negative | center)?
:margin-bottom=dimension?
:margin-left=dimension?
:margin-right=dimension?
:margin-top=dimension?
:textAlignment=(left | right | center | decimal | mixed)?
:decimal-offset=dimension?
=> EMPTY
borderProperties => border_style+
any[border_style]
:borderStyleType=(none | solid | dashed | thick | thin | double)?
:color=color?
=> EMPTY
printingProperties
:printAllLayers=bool?
:rescaleLongTableToFitPage=bool?
:rescaleWideTableToFitPage=bool?
:windowOrphanLines=int?
:continuationText?
:continuationTextAtBottom=bool?
:continuationTextAtTop=bool?

        :printEachLayerOnSeparatePage=bool?
     => EMPTY

The name attribute appears only in standalone .stt files (see Section 3.1 [SPSS TableLook STT Format], page 84).

3 SPSS TableLook File Formats

SPSS has a concept called a TableLook to control the styling of pivot tables in output. SPSS 15 and earlier used .tlo files with a special binary format to save TableLooks to disk; SPSS 16 and later use .stt files in an XML format to save them. Both formats expose roughly the same features, although the older .tlo format does have some features that .stt does not.

This chapter describes both formats.

3.1 The .stt Format

The .stt file format is an XML file that contains a subset of the SPV structure member format (see Section 2.1 [SPV Structure Member Format], page 29). Its root element is a tableProperties element (see Section 2.4.14 [SPV Detail Legacy Properties], page 81).

3.2 The .tlo Format

A .tlo file has a custom binary format. This section describes it using the syntax used previously for SPV binary members (see Section 2.2 [SPV Light Detail Member Format], page 40). There is one new convention: TLO files express colors as int32 values in which the low 8 bits are the red component, the next 8 bits are green, and next 8 bits are blue, and the high bits are zeros.

TLO files support various features that SPV files do not. PSPP implements the SPV feature set, so it mostly ignores the added TLO features. The details of this mapping are explained below.

At the top level, a TLO file consists of five sections. The first four are always present and the last one is optional:

     TableLook =>
        PTTableLook[tl]
        PVSeparatorStyle[ss]
        PVCellStyle[cs]
        PVTextStyle[ts]
        V2Styles?

Each section is described below.

3.2.1 PTTableLook
     PTTableLook =>
        ff ff 00 00 "PTTableLook" (00|02)[version]
        int16[flags]
        00 00
        bool[nested-row-labels] 00
        bool[footnote-marker-subscripts] 00
        i54 i18

In PTTableLook, version is 00 or 02. The only difference is that version 00 lacks V2Styles (see Section 3.2.4 [V2Styles in SPSS TLO Files], page 87) and that version 02 includes it. Both TLO versions are seen in the wild.

flags is a bit-mapped field. Its bits have the following meanings:

0x2 If set to 1, hide empty rows and columns; otherwise, show them.

0x4 If set to 1, use numeric footnote markers; otherwise, use alphabetic footnote markers.

0x8 If set to 1, print all layers; otherwise, print only the current layer.

0x10 If set to 1, scale the table to fit the page width; otherwise, break it horizontally if necessary.

0x20 If set to 1, scale the table to fit the page length; otherwise, break it vertically if necessary.

0x40 If set to 1, print each layer on a separate page (only if all layers are being printed); otherwise, paginate layers naturally.

0x80 If set to 1, print a continuation string at the top of a table that is split between pages.

0x100 If set to 1, print a continuation string at the bottom of a table that is split between pages.

When nested-row-labels is 1, row dimension labels appear nested; otherwise, they are put into the upper-left corner of the pivot table.

When footnote-marker-subscripts is 1, footnote markers are shown as subscripts; otherwise, they are shown as superscripts.

3.2.2 PVSeparatorStyle
     PVSeparatorStyle =>
        ff ff 00 00 "PVSeparatorStyle" 00
        Separator*4[sep1]
        03 80 00
        Separator*4[sep2]
     Separator =>
        case(
            00 00
          | 01 00 int32[color] int16[style] int16[width]
        )[type]

PVSeparatorStyle contains eight Separators, in two groups. Each Separator represents a border between pivot table elements. TLO and SPV files have the same concepts for borders. See Section 2.2.5 [SPV Light Member Borders], page 44, for the treatment of borders in SPV files.

A Separator’s type is 00 if the border is not drawn, 01 otherwise. For a border that is drawn, color is the color that it is drawn in. style and width have the following meanings:

style = 0 and 0 ≤ width ≤ 3

An increasingly thick single line. SPV files only have three line thicknesses. PSPP treats width 0 as a thin line, width 1 as a solid (normal width) line, and width 2 or 3 as a thick line.

style = 1 and 0 ≤ width ≤ 1

A doubled line, composed of normal-width (0) or thick (1) lines. SPV files only have “normal” width double lines, so PSPP maps both variants the same way.

style = 2 A dashed line.

The first group, sep1, represents the following borders within the pivot table, by index:

0. Horizontal dimension rows

1. Vertical dimension rows

2. Horizontal category rows

3. Vertical category rows

The second group, sep2, represents the following borders within the pivot table, by index:

0. Horizontal dimension columns

1. Vertical dimension columns

2. Horizontal category columns

3. Vertical category columns

3.2.3 PVCellStyle and PVTextStyle
     PVCellStyle =>
        ff ff 00 00 "PVCellStyle"
        AreaColor[title-color]
     PVTextStyle =>
        ff ff 00 00 "PVTextStyle" 00
        AreaStyle[title-style] MostAreas*7[most-areas]
     MostAreas =>
        06 80
        AreaColor[color] 08 80 00 AreaStyle[style]

These sections hold the styling and coloring for each of the 8 areas in a pivot table. They are conceptually similar to the area style information in SPV light members (see Section 2.2.4 [SPV Light Member Areas], page 44).

The styling and coloring for the title area is split between PVCellStyle and PVTextStyle: the former holds title-color, the latter holds title-style. The style for the remaining 7 areas is in most-areas in PVTextStyle, in the following order: layers, corner, row labels, column labels, data, caption, and footer.

     AreaColor =>
        00 01 00 int32[color10] int32[color0] byte[shading] 00

AreaColor represents the background color of an area. TLO files, but not SPV files, describe backgrounds that are a shaded combination of two colors: shading of 0 is pure color0, shading of 10 is pure color10, and value in between mix pixels of the two different colors in linear degree. PSPP does not implement shading, so for 1 ≤ shading ≤ 9 it interpolates RGB values between colors to arrive at an intermediate shade.

     AreaStyle =>

int16[valign] int16[halign] int16[decimal-offset]
int16[left-margin] int16[right-margin] int16[top-margin] int16[bottom-margin]
00 00 01 00
int32[font-size] int16[stretch]
00*2
int32[rotation-angle]
00*4
int16[weight]
00*2
bool[italic] bool[underline] bool[strikethrough]
int32[rtf-charset-number]
byte[x]
byte[font-name-len] byte*[font-name-len][font-name]
int32[text-color]
00*2

AreaStyle represents style properties of an area.

valign is 0 for top alignment, 1 for bottom alginment, 2 for center.

halign is 0 for left alignment, 1 for right, 2 for center, 3 for mixed, 4 for decimal. For decimal alignment, decimal-offset is the offset of the decimal point in 20ths of a point.

left-margin, right-margin, top-margin, and bottom-margin are also measured in

20ths of a point.

font-size is negative 96ths of an inch, e.g. 9 point is -12 or 0xfffffff3.

stretch has something to do with font size or stretch. The usual value is 01 and values larger than that do weird things. A reader can safely ignore it.

rotation-angle is a font rotation angle. A reader can safely ignore it.

weight is 400 for a normal-weight font, 700 indicates bold. (This is a Windows API convention.)

italic and underline have the obvious meanings. So does strikethrough, which PSPP ignores.

rtf-charset-number is a character set number from RTF. A reader can safely ignore it.

The meaning of x is unknown. Values 12, 22, 31, and 32 have been observed.

The font-name is the name of a font, such as Arial. Only US-ASCII characters have been observed here.

text-color is the color of the text itself.

3.2.4 V2Styles
V2Styles =>
Separator*11[sep3]
byte[continuation-len] byte*[continuation-len][continuation]
int32[min-col-width] int32[max-col-width]
int32[min-row-height] int32[max-row-height]

This final, optional, part of the TLO file format contains some additional style information. It begins with sep3, which represents the following borders within the pivot table, by index:

0 Title.

1…4 Left, right, top, and bottom inner frame.

5…8 Left, right, top, and bottom outer frame.

9, 10 Left and top of data area.

When V2Styles is absent, the inner frame borders default to a solid line and the others listed above to no line.

continuation is the string that goes at the top or bottom of a table broken across pages. When V2Styles is absent, the default is (Cont.).

min-col-width is the minimum width that a column will be assigned automatically. max-col-width is the maximum width that a column will be assigned to accommodate a long column label. min-row-width and max-row-width are a similar range for the width of row labels. All of these measurements are in points. When V2Styles is absent, the

defaults are 36 for min-col-width and min-row-height, 72 for max-col-width, and 120
for max-row-height.

4 Encrypted File Wrappers

SPSS 21 and later can package multiple kinds of files inside an encrypted wrapper. The wrapper has a common format, regardless of the kind of the file that it contains.

Warning: The SPSS encryption wrapper is poorly designed. When the password is unknown, it is much cheaper and faster to decrypt a file encrypted this way than if a well designed alternative were used. If you must use this format, use a 10-byte randomly generated password.

4.1 Common Wrapper Format

An encrypted file wrapper begins with the following 36-byte header, where xxx identifies the type of file encapsulated: SAV for a system file, SPS for a syntax file, SPV for a viewer file. PSPP code for identifying these files just checks for the ENCRYPTED keyword at offset 8, but the other bytes are also fixed in practice:

     0000  1c 00 00 00 00 00 00 00  45 4e 43 52 59 50 54 45  |........ENCRYPTE|
     0010  44 xx xx xx 15 00 00 00  00 00 00 00 00 00 00 00  |Dxxx............|
     0020  00 00 00 00                                       |....|

Following the fixed header is essentially the regular contents of the encapsulated file in its usual format, with each 16-byte block encrypted with AES-256 in ECB mode.

To make the plaintext an even multiple of 16 bytes in length, the encryption process appends PKCS #7 padding, as specified in RFC 5652 section 6.3. Padding appends 1 to 16 bytes to the plaintext, in which each byte of padding is the number of padding bytes added. If the plaintext is, for example, 2 bytes short of a multiple of 16, the padding is 2 bytes with value 02; if the plaintext is a multiple of 16 bytes in length, the padding is 16 bytes with value 0x10.

The AES-256 key is derived from a password in the following way:

1. Start from the literal password typed by the user. Truncate it to at most 10 bytes, then append as many null bytes as necessary until there are exactly 32 bytes. Call this password.

2. Let constant be the following 73-byte constant:

         0000  00 00 00 01 35 27 13 cc  53 a7 78 89 87 53 22 11
         0010  d6 5b 31 58 dc fe 2e 7e  94 da 2f 00 cc 15 71 80
         0020  0a 6c 63 53 00 38 c3 38  ac 22 f3 63 62 0e ce 85
         0030  3f b8 07 4c 4e 2b 77 c7  21 f5 1a 80 1d 67 fb e1
         0040  e1 83 07 d8 0d 00 00 01  00

3. Compute CMAC-AES-256(password, constant). Call the 16-byte result cmac.

4. The 32-byte AES-256 key is cmac || cmac, that is, cmac repeated twice.

Example

Consider the password ‘pspp’. password is:

     0000  70 73 70 70 00 00 00 00  00 00 00 00 00 00 00 00  |pspp............|
     0010  00 00 00 00 00 00 00 00  00 00 00 00 00 00 00 00  |................|

cmac is:

     0000  3e da 09 8e 66 04 d4 fd  f9 63 0c 2c a8 6f b0 45

The AES-256 key is:

     0000  3e da 09 8e 66 04 d4 fd  f9 63 0c 2c a8 6f b0 45
     0010  3e da 09 8e 66 04 d4 fd  f9 63 0c 2c a8 6f b0 45

4.1.1 Checking Passwords

A program reading an encrypted file may wish to verify that the password it was given is the correct one. One way is to verify that the PKCS #7 padding at the end of the file is well formed. However, any plaintext that ends in byte 01 is well formed PKCS #7, meaning that about 1 in 256 keys will falsely pass this test. This might be acceptable for interactive use, but the false positive rate is too high for a brute-force search of the password space.

A better test requires some knowledge of the file format being wrapped, to obtain a “magic number” for the beginning of the file.

• The plaintext of system files begins with $FL2@(#) or $FL3@(#).

• Before encryption, a syntax file is prefixed with a line at the beginning of the form * Encoding: encoding., where encoding is the encoding used for the rest of the file, e.g. windows-1252. Thus, * Encoding may be used as a magic number for system files.

• The plaintext of viewer files begins with 50 4b 03 04 14 00 08 (50 4b is PK).

4.2 Password Encoding

SPSS also supports what it calls “encrypted passwords.” These are not encrypted. They are encoded with a simple, fixed scheme. An encoded password is always a multiple of 2 characters long, and never longer than 20 characters. The characters in an encoded password are always in the graphic ASCII range 33 through 126. Each successive pair of characters in the password encodes a single byte in the plaintext password.

Use the following algorithm to decode a pair of characters:

1. Let a be the ASCII code of the first character, and b be the ASCII code of the second character.

2. Let ah be the most significant 4 bits of a. Find the line in the table below that has ah on the left side. The right side of the line is a set of possible values for the most significant 4 bits of the decoded byte.

         2 ⇒ 2367
         3 ⇒ 0145
         47 ⇒ 89cd
         56 ⇒ abef

3. Let bh be the most significant 4 bits of b. Find the line in the second table below that has bh on the left side. The right side of the line is a set of possible values for the most significant 4 bits of the decoded byte. Together with the results of the previous step, only a single possibility is left.

         2 ⇒ 139b
         3 ⇒ 028a
         47 ⇒ 46ce
         56 ⇒ 57df

4. Let al be the least significant 4 bits of a. Find the line in the table below that has al on the left side. The right side of the line is a set of possible values for the least significant 4 bits of the decoded byte.

03cf ⇒ 0145
12de ⇒ 2367
478b ⇒ 89cd
569a ⇒ abef

5. Let bl be the least significant 4 bits of b. Find the line in the table below that has bl on the left side. The right side of the line is a set of possible values for the least significant 4 bits of the decoded byte. Together with the results of the previous step, only a single possibility is left.

03cf ⇒ 028a
12de ⇒ 139b
478b ⇒ 46ce
569a ⇒ 57df

Example

Consider the encoded character pair ‘-|’. a is 0x2d and b is 0x7c, so ah is 2, bh is 7, al is 0xd, and bl is 0xc. ah means that the most significant four bits of the decoded character is 2, 3, 6, or 7, and bh means that they are 4, 6, 0xc, or 0xe. The single possibility in common is 6, so the most significant four bits are 6. Similarly, al means that the least significant four bits are 2, 3, 6, or 7, and bl means they are 0, 2, 8, or 0xa, so the least significant four bits are 2. The decoded character is therefore 0x62, the letter ‘b’.

5 Portable File Format

These days, most computers use the same internal data formats for integer and floatingpoint data, if one ignores little differences like big- versus little-endian byte ordering. However, occasionally it is necessary to exchange data between systems with incompatible data formats. This is what portable files are designed to do.

The portable file format is mostly obsolete. System files (see Chapter 1 [System File Format], page 2) are a better alternative.

Please note: This information is gleaned from examination of ASCII-formatted portable files only, so some of it may be incorrect for portable files formatted in EBCDIC or other character sets.

5.1 Portable File Characters

Portable files are arranged as a series of lines of 80 characters each. Each line is terminated by a carriage-return, line-feed sequence (“new-lines”). New-lines are only used to avoid line length limits imposed by some OSes; they are not meaningful.

Most lines in portable files are exactly 80 characters long. The only exception is a line that ends in one or more spaces, in which the spaces may optionally be omitted. Thus, a portable file reader must act as though a line shorter than 80 characters is padded to that length with spaces.

The file must be terminated with a ‘Z’ character. In addition, if the final line in the file does not have exactly 80 characters, then it is padded on the right with ‘Z’ characters. (The file contents may be in any character set; the file contains a description of its own character set, as explained in the next section. Therefore, the ‘Z’ character is not necessarily an ASCII ‘Z’.)

For the rest of the description of the portable file format, new-lines and the trailing ‘Z’s will be ignored, as if they did not exist, because they are not an important part of understanding the file contents.

5.2 Portable File Structure

Every portable file consists of the following records, in sequence:

• File header.

• Version and date info.

• Product identification.

• Author identification (optional).

• Subproduct identification (optional).

• Variable count.

• Case weight variable (optional).

• Variables. Each variable record may optionally be followed by a missing value record and a variable label record.

• Value labels (optional).

• Documents (optional).

• Data.

Most records are identified by a single-character tag code. The file header and version info record do not have a tag.

Other than these single-character codes, there are three types of fields in a portable file: floating-point, integer, and string. Floating-point fields have the following format:

• Zero or more leading spaces.

• Optional asterisk (‘*’), which indicates a missing value. The asterisk must be followed by a single character, generally a period (‘.’), but it appears that other characters may also be possible. This completes the specification of a missing value.

• Optional minus sign (‘-’) to indicate a negative number.

• A whole number, consisting of one or more base-30 digits: ‘0’ through ‘9’ plus capital letters ‘A’ through ‘T’.

• Optional fraction, consisting of a radix point (‘.’) followed by one or more base-30 digits.

• Optional exponent, consisting of a plus or minus sign (‘+’ or ‘-’) followed by one or more base-30 digits.

• A forward slash (‘/’).

Integer fields take a form identical to floating-point fields, but they may not contain a fraction.

String fields take the form of an integer field having value n, followed by exactly n characters, which are the string content.

5.3 Portable File Header

Every portable file begins with a 464-byte header, consisting of a 200-byte collection of vanity splash strings, followed by a 256-byte character set translation table, followed by an 8-byte tag string.

The 200-byte segment is divided into five 40-byte sections, each of which represents the string charset SPSS PORT FILE in a different character set encoding, where charset is the name of the character set used in the file, e.g. ASCII or EBCDIC. Each string is padded on the right with spaces in its respective character set.

It appears that these strings exist only to inform those who might view the file on a screen, and that they are not parsed by SPSS products. Thus, they can be safely ignored. For those interested, the strings are supposed to be in the following character sets, in the specified order: EBCDIC, 7-bit ASCII, CDC 6-bit ASCII, 6-bit ASCII, Honeywell 6-bit ASCII.

The 256-byte segment describes a mapping from the character set used in the portable file to an arbitrary character set having characters at the following positions:

0–60 Control characters. Not important enough to describe in full here.

61–63 Reserved.

64–73

Digits ‘0’ through ‘9’.

74–99

Capital letters ‘A’ through ‘Z’.

100–125

Lowercase letters ‘a’ through ‘z’.

126

Space.

127–130

Symbols .<(+

131

Solid vertical pipe.

132–142

          Symbols &[]!$*);^-/

143

Broken vertical pipe.

144–150

Symbols ,%_>?`:

151

British pound symbol.

152–155

Symbols @'=".

156

Less than or equal symbol.

157

Empty box.

158

Plus or minus.

159

Filled box.

160

Degree symbol.

161

Dagger.

162 Symbol ‘~’.

163

En dash.

164

Lower left corner box draw.

165

Upper left corner box draw.

166

Greater than or equal symbol.

167–176

Superscript ‘0’ through ‘9’.

177

Lower right corner box draw.

178

Upper right corner box draw.

179

Not equal symbol.

180

Em dash.

181

Superscript ‘(’.

182

Superscript ‘)’.

183

Horizontal dagger (?).

184–186

Symbols ‘{}\’.

187

Cents symbol.

188

Centered dot, or bullet.

189–255

Reserved. Symbols that are not defined in a particular character set are set to the same value as

symbol 64; i.e., to ‘0’.

The 8-byte tag string consists of the exact characters SPSSPORT in the portable file’s character set, which can be used to verify that the file is indeed a portable file.

5.4 Version and Date Info Record

This record does not have a tag code. It has the following structure:

• A single character identifying the file format version. The letter A represents version 0, and so on.

• An 8-character string field giving the file creation date in the format YYYYMMDD.

• A 6-character string field giving the file creation time in the format HHMMSS.

5.5 Identification Records

The product identification record has tag code ‘1’. It consists of a single string field giving the name of the product that wrote the portable file.

The author identification record has tag code ‘2’. It is optional. If present, it consists of a single string field giving the name of the person who caused the portable file to be written.

The subproduct identification record has tag code ‘3’. It is optional. If present, it consists of a single string field giving additional information on the product that wrote the portable file.

5.6 Variable Count Record

The variable count record has tag code ‘4’. It consists of a single integer field giving the number of variables in the file dictionary.

5.7 Precision Record

The precision record has tag code ‘5’. It consists of a single integer field specifying the maximum number of base-30 digits used in data in the file.

5.8 Case Weight Variable Record

The case weight variable record is optional. If it is present, it indicates the variable used for weighting cases; if it is absent, cases are unweighted. It has tag code ‘6’. It consists of a single string field that names the weighting variable.

5.9 Variable Records

Each variable record represents a single variable. Variable records have tag code ‘7’. They have the following structure:

• Width (integer). This is 0 for a numeric variable, and a number between 1 and 255 for a string variable.

• Name (string). 1–8 characters long. Must be in all capitals.

A few portable files that contain duplicate variable names have been spotted in the wild. PSPP handles these by renaming the duplicates with numeric extensions: var_1, var_2, and so on.

• Print format. This is a set of three integer fields:

− Format type (see Section 1.3 [Variable Record], page 6).

− Format width. 1–40.

− Number of decimal places. 1–40.

A few portable files with invalid format types or formats that are not of the appropriate width for their variables have been spotted in the wild. PSPP assigns a default F or A format to a variable with an invalid format.

• Write format. Same structure as the print format described above.

Each variable record can optionally be followed by a missing value record, which has tag code ‘8’. A missing value record has one field, the missing value itself (a floating-point or string, as appropriate). Up to three of these missing value records can be used.

There is also a record for missing value ranges, which has tag code ‘B’. It is followed by two fields representing the range, which are floating-point or string as appropriate. If a missing value range is present, it may be followed by a single missing value record.

Tag codes ‘9’ and ‘A’ represent LO THRU x and x THRU HI ranges, respectively. Each is followed by a single field representing x. If one of the ranges is present, it may be followed by a single missing value record.

In addition, each variable record can optionally be followed by a variable label record, which has tag code ‘C’. A variable label record has one field, the variable label itself (string).

5.10 Value Label Records

Value label records have tag code ‘D’. They have the following format:

• Variable count (integer).

• List of variables (strings). The variable count specifies the number in the list. Variables are specified by their names. All variables must be of the same type (numeric or string), but string variables do not necessarily have the same width.

• Label count (integer).

• List of (value, label) tuples. The label count specifies the number of tuples. Each tuple consists of a value, which is numeric or string as appropriate to the variables, followed by a label (string).

A few portable files that specify duplicate value labels, that is, two different labels for a single value of a single variable, have been spotted in the wild. PSPP uses the last value label specified in these cases.

5.11 Document Record

One document record may optionally follow the value label record. The document record consists of tag code ‘E’, following by the number of document lines as an integer, followed by that number of strings, each of which represents one document line. Document lines must be 80 bytes long or shorter.

5.12 Portable File Data

The data record has tag code ‘F’. There is only one tag for all the data; thus, all the data must follow the dictionary. The data is terminated by the end-of-file marker ‘Z’, which is not valid as the beginning of a data element.

Data elements are output in the same order as the variable records describing them. String variables are output as string fields, and numeric variables are output as floatingpoint fields.

6 SPSS/PC+ System File Format

SPSS/PC+, first released in 1984, was a simplified version of SPSS for IBM PC and compatible computers. It used a data file format related to the one described in the previous chapter, but simplified and incompatible. The SPSS/PC+ software became obsolete in the 1990s, so files in this format are rarely encountered today. Nevertheless, for completeness, and because it is not very difficult, it seems worthwhile to support at least reading these files. This chapter documents this format, based on examination of a corpus of about 60 files from a variety of sources.

System files use four data types: 8-bit characters, 16-bit unsigned integers, 32-bit unsigned integers, and 64-bit floating points, called here char, uint16, uint32, and flt64, respectively. Data is not necessarily aligned on a word or double-word boundary.

SPSS/PC+ ran only on IBM PC and compatible computers. Therefore, values in these files are always in little-endian byte order. Floating-point numbers are always in IEEE 754 format.

SPSS/PC+ system files represent the system-missing value as -1.66e308, or f5 1e 26 02 8a 8c ed ff expressed as hexadecimal. (This is an unusual choice: it is close to, but not equal to, the largest negative 64-bit IEEE 754, which is about -1.8e308.)

Text in SPSS/PC+ system file is encoded in ASCII-based 8-bit MS DOS codepages. The corpus used for investigating the format were all ASCII-only.

An SPSS/PC+ system file begins with the following 256-byte directory:

     uint32              two;
     uint32              zero;
     struct {
         uint32          ofs;
         uint32          len;
     } records[15];
     char                filename[128];
uint32 two;
uint32 zero;

Always set to 2 and 0, respectively.

These fields could be used as a signature for the file format, but the product field in record 0 seems more likely to be unique (see Section 6.1 [Record 0 Main Header Record], page 100).

struct { ... } records[15];

Each of the elements in this array identifies a record in the system file. The ofs is a byte offset, from the beginning of the file, that identifies the start of the record. len specifies the length of the record, in bytes. Many records are optional or not used. If a record is not present, ofs and len for that record are both are zero.

char filename[128];

In most files in the corpus, this field is entirely filled with spaces. In one file, it contains a file name, followed by a null bytes, followed by spaces to fill the remainder of the field. The meaning is unknown.

The following sections describe the contents of each record, identified by the index into the records array.

6.1 Record 0: Main Header Record

All files in the corpus have this record at offset 0x100 with length 0xb0 (but readers should find this record, like the others, via the records table in the directory). Its format is:

uint16              one0;
char                product[62];
flt64               sysmis;
uint32              zero0;
uint32              zero1;
uint16              one1;
uint16              compressed;
uint16              nominal_case_size;
uint16              n_cases0;
uint16              weight_index;
uint16              zero2;
uint16              n_cases1;
uint16              zero3;
char                creation_date[8];
char                creation_time[8];
char                label[64];
uint16 one0;
uint16 one1;

Always set to 1.

uint32 zero0;
uint32 zero1;
uint16 zero2;
uint16 zero3;

Always set to 0.

It seems likely that one of these variables is set to 1 if weighting is enabled, but none of the files in the corpus is weighted.

char product[62];

Name of the program that created the file. Only the following unique values have been observed, in each case padded on the right with spaces:

          DESPSS/PC+ System File Written by Data Entry II
          PCSPSS SYSTEM FILE.  IBM PC DOS, SPSS/PC+
          PCSPSS SYSTEM FILE.  IBM PC DOS, SPSS/PC+ V3.0
          PCSPSS SYSTEM FILE.  IBM PC DOS, SPSS for Windows

Thus, it is reasonable to use the presence of the string ‘SPSS’ at offset 0x104 as a simple test for an SPSS/PC+ data file.

flt64 sysmis;

The system-missing value, as described previously (see Chapter 6 [SPSS/PC+ System File Format], page 99).

uint16 compressed;

Set to 0 if the data in the file is not compressed, 1 if the data is compressed with simple bytecode compression.

uint16 nominal_case_size;

Number of data elements per case. This is the number of variables, except that long string variables add extra data elements (one for every 8 bytes after the first 8). String variables in SPSS/PC+ system files are limited to 255 bytes.

uint16 n_cases0;
uint16 n_cases1;

The number of cases in the data record. Both values are the same. Some files in the corpus contain data for the number of cases noted here, followed by garbage that somewhat resembles data.

uint16 weight_index;

0, if the file is unweighted, otherwise a 1-based index into the data record of the weighting variable, e.g. 4 for the first variable after the 3 system-defined variables.

char creation_date[8];

The date that the file was created, in ‘mm/dd/yy’ format. Single-digit days and months are not prefixed by zeros. The string is padded with spaces on right or left or both, e.g. ‘_2/4/93_’, ‘10/5/87_’, and ‘_1/11/88’ (with ‘_’ standing in for a space) are all actual examples from the corpus.

char creation_time[8];

The time that the file was created, in ‘HH:MM:SS’ format. Single-digit hours are padded on a left with a space. Minutes and seconds are always written as two digits.

char file_label[64];

File label declared by the user, if any (see Section “FILE LABEL” in PSPP Users Guide). Padded on the right with spaces.

6.2 Record 1: Variables Record

The variables record most commonly starts at offset 0x1b0, but it can be placed elsewhere. The record contains instances of the following 32-byte structure:

uint32              value_label_start;
uint32              value_label_end;
uint32              var_label_ofs;
uint32              format;
char                name[8];
union {
flt64           f;
char            s[8];
} missing;

The number of instances is the nominal_case_size specified in the main header record. There is one instance for each numeric variable and each string variable with width 8 bytes or less. String variables wider than 8 bytes have one instance for each 8 bytes, rounding up. The first instance for a long string specifies the variable’s correct dictionary information. Subsequent instances for a long string are generally filled with all-zero bytes, although the missing field contains the numeric system-missing value, and some writers also fill in var_ label_ofs, format, and name, sometimes filling the latter with the numeric system-missing value rather than a text string. Regardless of the values used, readers should ignore the contents of these additional instances for long strings.

uint32 value_label_start;
uint32 value_label_end;

For a variable with value labels, these specify offsets into the label record of the start and end of this variable’s value labels, respectively. See Section 6.3 [Record 2 Labels Record], page 103, for more information. For a variable without any value labels, these are both zero. A long string variable may not have value labels.

uint32 var_label_ofs;

For a variable with a variable label, this specifies an offset into the label record. See Section 6.3 [Record 2 Labels Record], page 103, for more information. For a variable without a variable label, this is zero.

uint32 format;

The variable’s output format, in the same format used in system files. See [System File Output Formats], page 8, for details. SPSS/PC+ system files only use format types 5 (F, for numeric variables) and 1 (A, for string variables).

char name[8];

The variable’s name, padded on the right with spaces.

union { ... } missing;

A user-missing value. For numeric variables, missing.f is the variable’s usermissing value. For string variables, missing.s is a string missing value. A variable without a user-missing value is indicated with missing.f set to the system-missing value, even for string variables (!). A Long string variable may not have a missing value.

In addition to the user-defined variables, every SPSS/PC+ system file contains, as its first three variables, the following system-defined variables, in the following order. The system-defined variables have no variable label, value labels, or missing values.

$CASENUM A numeric variable with format F8.0. Most of the time this is a sequence number, starting with 1 for the first case and counting up for each subsequent case. Some files skip over values, which probably reflects cases that were deleted.

$DATE A string variable with format A8. Same format (including varying padding) as the creation_date field in the main header record (see Section 6.1 [Record 0 Main Header Record], page 100). The actual date can differ from creation_ date and from record to record. This may reflect when individual cases were added or updated.

$WEIGHT A numeric variable with format F8.2. This represents the case’s weight; SPSS/PC+ files do not have a user-defined weighting variable. If weighting has not been enabled, every case has value 1.0.

6.3 Record 2: Labels Record

The labels record holds value labels and variable labels. Unlike the other records, it is not meant to be read directly and sequentially. Instead, this record must be interpreted one piece at a time, by following pointers from the variables record.

   The value_label_start, value_label_end, and var_label_ofs fields in a variable

record are all offsets relative to the beginning of the labels record, with an additional 7byte offset. That is, if the labels record starts at byte offset labels_ofs and a variable has a given var_label_ofs, then the variable label begins at byte offset labels_ofs +

var_label_ofs + 7 in the file.

A variable label, starting at the offset indicated by var_label_ofs, consists of a one-byte length followed by the specified number of bytes of the variable label string, like this:

     uint8               length;
     char                s[length];

A set of value labels, extending from value_label_start to value_label_end (exclusive), consists of a numeric or string value followed by a string in the format just described. String values are padded on the right with spaces to fill the 8-byte field, like this:

     union {
         flt64           f;
         char            s[8];
     } value;
     uint8               length;
     char                s[length];

The labels record begins with a pair of uint32 values. The first of these is always 3. The second is between 8 and 16 less than the number of bytes in the record. Neither value is important for interpreting the file.

6.4 Record 3: Data Record

The format of the data record varies depending on the value of compressed in the file header record:

0: no compression Data is arranged as a series of 8-byte elements, one per variable instance variable in the variable record (see Section 6.2 [Record 1 Variables Record], page 101). Numeric values are given in flt64 format; string values are literal characters string, padded on the right with spaces when necessary to fill out 8-byte units.

1: bytecode compression The first 8 bytes of the data record is divided into a series of 1-byte command codes. These codes have meanings as described below:

0 The system-missing value.

1 A numeric or string value that is not compressible. The value is stored in the 8 bytes following the current block of command bytes. If this value appears twice in a block of command bytes, then it indicates the second group of 8 bytes following the command bytes, and so on.

2 through 255 A number with value code - 100, where code is the value of the compression code. For example, code 105 indicates a numeric variable of value 5.

The end of the 8-byte group of bytecodes is followed by any 8-byte blocks of non-compressible values indicated by code 1. After that follows another 8-byte group of bytecodes, then those bytecodes’ non-compressible values. The pattern repeats up to the number of cases specified by the main header record have been seen. The corpus does not contain any files with command codes 2 through 95, so it is possible that some of these codes are used for special purposes.

Cases of data often, but not always, fill the entire data record. Readers should stop reading after the number of cases specified in the main header record. Otherwise, readers may try to interpret garbage following the data as additional cases.

6.5 Records 4 and 5: Data Entry

Records 4 and 5 appear to be related to SPSS/PC+ Data Entry.

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.